### v1

# supplement_jasa

Abstract
Building artificially intelligent geospatial systems require rapid delivery of spatial data analysis at massive scales with minimal human intervention. Depending upon their intended use, data analysis may also entail model assessment and uncertainty quantification. This article devises transfer learning frameworks for deployment in artificially intelligent systems, where a massive data set is split into smaller data sets that stream into the analytical framework to propagate learning and assimilate inference for the entire data set. Specifically, we introduce Bayesian predictive stacking for multivariate spatial data and demonstrate its effectiveness in rapidly analyzing massive data sets. Furthermore, we make inference feasible in a reasonable amount of time, and without excessively demanding hardware settings. We illustrate the effectiveness of this approach in extensive simulation experiments and subsequently analyze massive data sets in climate science on sea surface temperatures and on vegetation index.
Keywords— Bayesian predictive stacking; Climate science; Gaussian process; Geospatial systems; Matrix-variate spatial models; Transfer learning.
1 Introduction
Geospatial artificial intelligence (GeoAI) is a rapidly evolving discipline at the interface of machine learning and spatial data science that attempts to harness the analytical capabilities of Artificial Intelligence (AI) to analyze massive amounts of geographic data and conduct data driven scientific discoveries in diverse domains within the environmental and physical sciences. The area, while still fledgling, presents opportunities for statisticians to devise essential data analytic tools that should comprise an artificially intelligent “geospatial” system. Constructing comprehensive AI-based geospatial systems involve expertise in computer science relating to efficiently designed Geographic Information Systems (GIS) databases, conventional machine learning tools based upon algorithms for spatial interpolation, and statistical or probabilistic learning tools to offer scientific inference. This manuscript concerns itself primarily with statistical learning tools for artificially intelligent GeoAI systems.
What, if at all, should be the role of formal statistical inference in GeoAI? Formal inference for spatial and spatial-temporal random fields enjoy a venerable intellectual presence spanning several decades of theoretical developments within classical and Bayesian paradigms (see, e.g., Cressie,, 1993; Stein,, 1999; Gelfand et al.,, 2010; Cressie and Wikle,, 2011; Banerjee et al.,, 2015). With increasing complexity in spatially oriented datasets, statisticians have built richly structured hierarchical models in their pursuit of “full” probabilistic inference. The term “full” loosely refers to approaches that attempt to estimate all unknowns in the model including all parameters (irrespective of how well they are identified by the data), random effects or partial realizations of the process, and predictive random variables for probabilistic interpolation at arbitrary locations (and time points in spatial-temporal data).
Statisticians would, we believe, broadly agree that GeoAI offers sufficient intellectual space to accommodate rigorous probabilistic learning as a crucial constituent in artificially intelligent data analysis systems. Spatial modeling relies upon Gaussian processes (GPs) to model dependence and achieves superior predictive inference. While GPs afford flexibility and are a conspicuous choice in spatial modeling, their covariance kernels do not necessarily yield computationally exploitable structures for the covariance matrices and the computational requirements for full inference become onerous and impracticable for massive datasets.
This presents a conundrum for statisticians: should we strive to completely retain the probabilistic rigors of theoretical inference, which may be computationally burdensome, or even impracticable, given the massive amounts of data, or should we concede uncertainty quantification and adopt algorithm-driven data analysis that would easily scale massive data sets? Spatial data analysis at unprecedented scales requires some concessions from conventional decision-theoretic paradigms as has been reflected, for example, in a recent comprehensive case study by Zhou et al., (2022), who synthesize statistical models with machine learning algorithms to measure temporal trends and spatial distribution of housing vitality by exploiting information from multiple sources (also see the discussion by Banerjee,, 2022, for conventional Bayesian modeling perspectives and its challenges). Other efforts in melding machine learning methods with formal spatial statistics include deep learning for spatial data (Zammit-Mangion et al.,, 2022; Wikle and Zammit-Mangion,, 2023, with the latter offering a comprehensive review of spatial machine learning) and, in a very related domain, for deep GP emulation in computer experiments (Sauer et al., 2023b, ), and spatial random forests (Georganos et al.,, 2021; Talebi et al.,, 2022; Saha et al.,, 2023).
Even a cursory review reveals a significant literature on statistical methods for massive spatial datasets, which is too vast to be summarized here (see, e.g., Banerjee,, 2017; Heaton et al.,, 2017, and references therein). Within the hierarchical setup $[data\mid process]\times[process\mid parameters]\times[parameters]$, inference proceeds from spatial processes that scale massive data sets. Examples range from reduced-rank processes or subsets of regression approaches (Quiñonero-Candela and Rasmussen,, 2005; Cressie and Johannesson,, 2008; Banerjee et al.,, 2008; Wikle,, 2010, and references therein), multi-resolution approaches (Nychka et al.,, 2015; Katzfuss,, 2017), and graph-based models Vecchia, (1988); Datta et al., (2016); Katzfuss and Guinness, (2021); Dey et al., (2022); Sauer et al., 2023a . Full inference typically requires efficient Markov chain Monte Carlo (mcmc) (Finley et al.,, 2019), variational approximations (Ren et al.,, 2011; Wu et al.,, 2022; Cao et al.,, 2023) and a significant body of literature on Gaussian Markov random field approximations (Rue et al.,, 2009; Lindgren et al.,, 2011, and references therein) in conjunction with integrated nested Laplace approximations (inla) for computing the marginal distributions of the process at given locations.
The aforementioned methods focus upon richness of statistical inference, but almost invariably involve a significant amount of human intervention to analyze spatial data. Even the simplest of geostatistical datasets require exploratory data analysis to learn about aspects of the underlying process that are weakly identified by the data (Zhang,, 2004; Tang et al.,, 2021). Building a GeoAI system, on the other hand, will require minimizing human intervention in offering a robust framework for spatial data analysis. This presents enough challenges that preclude a comprehensive solution in its entirety within the scope of a single manuscript. Nevertheless, we devise a spatial data analytic framework that holds significant promise for GeoAI. The premise of this approach relies upon two basic tenets: (i) model-based statistical inference for underlying spatial processes (including multivariate processes) in a robust and largely automated manner with minimal human input; and (ii) achieving such inference for truly massive amounts of data without resorting to iterative algorithms that may require significant human intervention to diagnose convergence (such as in mcmc).
We wish to retain the benefits of Bayesian hierarchical models, convincingly documented in the aforementioned manuscripts, with regard to spatial interpolation of the underlying stochastic process. In particular, we seek predictive inference over the uncountable collection of points in the spatial domain while propagating uncertainty about the model’s estimation in such inference. However, some compromises with regard to richness of models are inevitable from the standpoint of GeoAI. At the same time, statistical distribution theory has much to offer and we harness analytical closed-form distribution theory within a family of matrix-variate distributions to deliver inference. This is possible only by “fixing” certain spatial correlation kernel parameters that are weakly identified and hinder automated inference. Our approach relies upon carrying out fast and exact inference based upon the matrix-variate distributions and then assimilating the inference using Bayesian predictive stacking or bps (Wolpert,, 1992; Yao et al.,, 2018). Stacking (Breiman,, 1996) is primarily related to predictions by minimizing the generalization error rate of one or more models.
To analyze massive amounts of data to be accessed in a GeoAI system, we devise Bayesian transfer learning over a streaming set of spatial datasets. We envisage a setting where the data stream into the GeoAI system in the form of $K$ subsets, where each subset is composed of $\approx n/K$ locations with $K$ being a function of computational resources available to GeoAI. We analyze and transfer learning from one dataset to the next and assimilate inference from these subsets to the full data set over $n$ locations. The concept of delegating Bayesian calculations to a group of independent datasets is intuitive and has been explored from diverse perspectives. Examples include Consensus Monte Carlo (cmc) (Scott et al.,, 2016) and its kernel-based adaptations (Rendell et al.,, 2021), Bayesian meta-analysis in clinical applications (Parmigiani,, 2002, see, e.g., Chapter 4 in) and in spatially-temporally structured data (Bell et al.,, 2005; Kang et al.,, 2011); and, more recently, approaches such as “meta-kriging” using geometric medians of posterior distributions (Srivastava et al.,, 2015; Minsker et al.,, 2017; Guhaniyogi and Banerjee,, 2018, 2019; Guhaniyogi et al.,, 2019). However, each of these approaches have, hitherto, relied upon MCMC or other iterative algorithms for estimating the weakly identifiable parameter. We, instead, propose to execute transfer learning using a double stacking approach. We first obtain stacked inference for each of the parameters within each subset and, then, assimilate inference across subsets by a second stacking algorithm. A flowchart is supplied in Figure 1 comprising three steps: (i) partition the data into smaller subsets; (ii) analyze each subset by stacking conjugate posteriors in closed form; and (iii) stack across data sets to deliver inference for the entire data.
[…]
Figure 1: Double Bayesian predictive stacking approach representation
The contents of the paper evolve in accordance with Figure 1. Section 2 outlines Bayesian transfer learning using double bps. Section 3 provides details on computational aspects. Section 4 presents simulation experiments designed for empirically evaluating double bps. Section 5 analyzes a very large climate data set with observed locations in the millions at global scales. Finally, Section 6 concludes the paper with some discussion.
2 Bayesian transfer learning
Transfer learning (TL) broadly refers to propagating knowledge from one task to accomplish a different task. We consider transferring inference from one subset to the next in a stream of subsets to analyze the entire dataset. Learning from spatial random fields is challenging due to inherent dependencies in the data and we seek to infer about a latent spatial process (uncountable collection of unobserved random variables) with limited data (partially observed finite realization of the process). We develop an automated approach for GeoAI.
2.1 Divide-and-Conquer Multivariate Bayesian inference
Let $Y_{n\times q}$ be an $n\times q$ random matrix that is endowed with a probability law from the matrix-normal distribution, $\operatorname{MN}(M,V,U)$, with probability density function
$p(Y\mid M,V,U)=\frac{\exp\left[-\frac{1}{2}\operatorname{tr}\left\{U^{-1}(Y-M)^{\mathrm{\scriptscriptstyle T}}V^{-1}(Y-M)\right\}\right]}{(2\pi)^{\frac{np}{2}}\mid U\mid^{\frac{n}{2}}\mid V\mid^{\frac{q}{2}}},$ (2.1)
where $\operatorname{tr}(\cdot)$ is the trace operator on a square matrix, $M$ is the mean matrix, and $V$ and $U$ are the $n\times n$ row-covariance and $q\times q$ column covariance matrices, respectively. We consider the matrix-variate Bayesian linear regression model
$\begin{split}Y&=X\beta+E_{Y}\;,\;E_{Y}\sim\operatorname{MN}(O,V,\Sigma)\;;\;\beta=M_{0}m_{0}+E_{\beta}\;,\;E_{\beta}\sim\operatorname{MN}(O,M_{0},\Sigma)\;,\end{split}$ (2.2)
where $Y$ is $n\times q$, $X$ is $n\times p$ comprising explanatory variables, $\beta$ is $p\times q$ consisting of regression slopes, $E_{Y}$ and $E_{\beta}$ are zero-centered random matrices with row covariances $V$ and $M_{0}$, respectively, and a shared column covariance matrix $\Sigma$. We assign an inverse-Wishart distribution $\Sigma\sim\operatorname{IW}(\Psi_{0},\nu_{0})$ and denote the joint density of $\beta$ and $\Sigma$ by $\operatorname{MNIW}(M_{0}m_{0},M_{0},\Psi_{0},\nu_{0})$, which affords a closed-form posterior for $\{\beta,\Sigma\}$ in the same family.
Let $\mathscr{D}=\{Y,X\}$ be the entire dataset, which is too massive to even be accessed, let alone be analyzed using (2.1), within the GeoAI system. Therefore, we envisage $K$ disjoint and exhaustive subsets $\mathscr{D}=\{\mathscr{D}_{1},\dots,\mathscr{D}_{K}\}$ streaming into GeoAI as a sequence. Each $\mathscr{D}_{k}=\{Y_{k},X_{k}\}$ consists of $n_{k}$ rows of $Y$ and $X$, where $n=\sum_{k=1}^{K}n_{k}$, $Y_{k}$ is $n_{k}\times q$ and $X_{k}$ is $n_{k}\times p$ for each $k=1,\dots,K$. We now fit (2.2) to each subset using $Y_{k}=X_{k}\beta+E_{k}$, $E_{k}\sim\operatorname{MN}(O,V_{k},\Sigma)$, where $V_{k}$ is the $n_{k}\times n_{k}$ row-covariance matrix corresponding to the rows in $Y_{k}$; the specification for $\{\beta,\Sigma\}$ remains as in (2.2). Starting with $\operatorname{MNIW}(\beta,\Sigma\,|\,M_{k}m_{k},M_{k},\Psi_{k},\nu_{k})$ at $k=0$ (the prior), we use Bayesian updating $p(\beta,\Sigma\mid\mathscr{D}_{1:k+1})\propto p(\beta,\Sigma\mid\mathscr{D}_{1:k})\times p(Y_{k+1}\mid X_{k+1}\beta,V_{k+1},\Sigma)$ to obtain $\beta,\Sigma\mid\mathscr{D}_{1:k+1}\sim\operatorname{MNIW}(M_{k+1}m_{k+1},M_{k+1},\Psi_{k+1},\nu_{k+1})$ with $M_{k+1}^{-1}=M_{k}^{-1}+X_{k+1}^{\mathrm{\scriptscriptstyle T}}V_{k+1}^{-1}X_{k+1}$, $m_{k+1}=m_{k}+X_{k+1}^{\mathrm{\scriptscriptstyle T}}V_{k+1}^{-1}Y_{k+1}$, $\nu_{k+1}=\nu_{k}+n_{k+1}$, and $\Psi_{k+1}=\Psi_{k}+Y_{k+1}^{\mathrm{\scriptscriptstyle T}}V_{k+1}^{-1}Y_{k+1}+m_{k}^{\mathrm{\scriptscriptstyle T}}M_{k}m_{k}-m_{k+1}^{\mathrm{\scriptscriptstyle T}}M_{k+1}m_{k+1}$. Upon termination at $k=K$, we exactly recover the posterior distribution $p(\beta,\Sigma\,|\,\mathscr{D})$. There is no need to interact between the subsets, and computational complexity is determined solely by the dimension of the subsets.
Spatial random fields immediately present a challenge. The above method delivers full inference without any loss of information only if $Y_{k}$’s are reasonably assumed to be independent (or exchangeable) across blocks. If each row of $Y$ corresponds to one of $n$ spatial locations so that $V$ is an $n\times n$ spatial correlation matrix, the above scheme requires that each $V_{k}$ is the spatial correlation matrix constructed from the $n_{k}$ spatial locations in $\mathscr{D}_{k}$. Independence across the $K$ blocks may yield reasonable inference, especially if it is possible to design the blocks such that the spatial correlation across blocks do not impact overall inference, but this will require significant human intervention or a separate AI system for designing such blocks. We seek to avoid this in GeoAI.
Instead, we devise a method for assimilating the learning from each of these blocks using predictive stacking. We exploit the fact that $V$ is indexed by a small number of parameters in a spatial correlation kernel. Fixing these parameters fixes $V$, and hence $V_{k}$ for each $k$, yielding closed-form posterior inference on $\beta$ and $\Sigma$ based upon the entire dataset as described above. Stacking combines these analytically accessible distributions using an optimal set of weights that are computed using a convex optimization algorithm. These weights are then used to reconstruct the posterior and predictive distributions for the spatial random field without imposing block independence. It is worth remarking here that existing “meta-kriging” and related “divide and conquer” approaches (e.g., Guhaniyogi and Banerjee,, 2018; Scott et al.,, 2016, and other references on “divide and conquer” methods provided in Section 1) analyze subsets of data using mcmc, which is expensive and not fully automated. A key distinction of the current manuscript is that we abandon all iterative estimation algorithms, let alone expensive mcmc, and focus on assimilating inference using closed forms of distributions. The matrix-variate formulation we pursue enhances flexibility and scope of applications.
What remains to be resolved is the issue of fixing the parameters in $V$. These parameters govern the strength of association across the spatial random field and possibly the smoothness of the field. Unfortunately, these parameters are weakly identified by the data and posterior learning struggles due to slower convergence of iterative algorithms. Also, exploratory spatial data science tools, such as variograms, for gleaning information about these parameters will also be less helpful as they may suggest different values for each of the variables indexed by the columns, while we insist on retaining a single parameter to exploit conjugate distribution theory. Hence, we collect the closed-form posterior distributions obtained for a collection of fixed values of the spatial parameters and, subsequently, average these posterior distributions.
2.2 Bayesian stacking of predictive densities
Bayesian predictive stacking (bps) assimilates models using a weighted distribution in the convex hull, $C=\left\{\sum_{j=1}^{J}w_{j}p(\cdot\mid\mathscr{D},\mathscr{M}_{j}):\sum_{j}w_{j}=1,w_{j}\geq 0\right\}$, of individual posterior distributions by maximizing the score (Gneiting and Raftery,, 2007; Yao et al.,, 2018) to fetch
$(w_{1},\ldots,w_{J})^{{\mathrm{\scriptscriptstyle T}}}=\arg\max_{w\in S_{1}^{J}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{j=1}^{J}w_{j}p\left(Y_{i}\mid\mathscr{D}_{-i},\mathscr{M}_{j}\right)\;,$ (2.3)
where $\mathscr{D}_{-i}$ is the dataset excluding the $i$-th observation, or block (indexed by a row) of observations in the matrix $Y$, and $\mathscr{M}=\left(\mathscr{M}_{1},\ldots,\mathscr{M}_{J}\right)$ are $J$ different models. For any given dataset $\mathscr{D}$, these $J$ different models correspond to fixed spatial correlation kernel parameters in $V$ and deliver closed-form analytic distributions. Solving (2.3) minimizes the Kullback-Leibler divergence from the true predictive distribution and is executed using convex optimization (Grant and Boyd,, 2008; Fu et al.,, 2020). While the true predictive distribution is unknown, we can use a leave-one-out (loo) estimate of the expected value of the score (see. e.g., Yao et al.,, 2018, for details), which requires fitting the model $n$ times as we exclude one row of the data $\mathscr{D}$ at a time so we prefer K-fold cross-validation as a more cost-effective method for generating predictions (Breiman,, 1996).
2.3 Accelerated learning for spatial random fields
We turn to adapting the aforementioned concepts into an automated learning framework for GeoAI. Let $\mathcal{S}=\left\{s_{1},\ldots,s_{n}\right\}\subset\mathcal{D}$ be a set of $n$ locations yielding observations on $q$ possibly correlated outcomes collected into a $q\times 1$ vector $y(s)=(y_{1}(s),\ldots,y_{q}(s))^{{\mathrm{\scriptscriptstyle T}}}$ for each $s\in\mathcal{S}$. We collect these measurements into the $n\times q$ matrix $Y=[y_{j}(s_{i})^{{\mathrm{\scriptscriptstyle T}}}]$ for $i=1,\ldots,n$ and $j=1,\ldots,q$. Let $X=[x(s_{i})^{{\mathrm{\scriptscriptstyle T}}}]$ be $n\times p$ with rows $x(s_{i})^{{\mathrm{\scriptscriptstyle T}}}$ consisting of $p<n$ explanatory variables at location $s_{i}\in\mathcal{S}$; we assume $X$ has rank $p$. We cast this into (2.2), but explicitly introduce a latent $q\times 1$ spatial process $\omega(s)$ as
$\begin{split}Y&=X\beta+\Omega+E\;,\quad E\,|\,\Sigma\sim\operatorname{MN}(O,(\alpha^{-1}-1)\mathbb{I}_{n},\Sigma)\;;\quad\Sigma\sim\operatorname{IW}(\Psi_{0},\nu_{0})\;;\\ \beta&=M_{0}m_{0}+E_{\beta}\;,\quad E_{\beta}\,|\,\Sigma\sim\operatorname{MN}(O,M_{0},\Sigma)\;;\quad\Omega\,|\,\Sigma\sim\operatorname{MN}(O,V,\Sigma)\;,\end{split}$ (2.4)
where $\Omega=[\omega(s_{i})^{{\mathrm{\scriptscriptstyle T}}}]$ is $n\times q$ with rows $\omega(s_{i})^{{\mathrm{\scriptscriptstyle T}}}$. To capture spatial dependence, $V$ is an $n\times n$ spatial correlation matrix with $(i,j)$-th element equaling the value of a positive definite spatial correlation function $\rho(s_{i},s_{j};\phi)$ indexed by parameter(s) $\phi$. To account for measurement errors in observations, as is customary in geostatistics, we introduce a discontinuity in the spatial correlation function and modify the elements to $\rho(s_{i},s_{j};\phi)+(\alpha^{-1}-1)\mathds{1}_{s_{i}=s_{j}}$, where $\alpha\in[0,1]$ represents the proportion of total variability attributed to the spatial process.
The covariance among the elements of $y(s)$ within $s$ is given by the elements of $\left(\alpha^{-1}-1\right)\Sigma$. Thus, $\Sigma$ is the within-location (nonspatial) dependence among the outcomes adjusted by a scale of $\left(\alpha^{-1}-1\right)$ to accommodate additional variation at local scales. The parameter $\alpha$ is the ratio of the “partial sill” to the “sill” in classical geostatistics. In the special case when $\Sigma=\sigma^{2}\mathbb{I}_{q}$, we interpret $\alpha=\sigma^{2}/\left(\sigma^{2}+\tau^{2}\right)$, where $\tau^{2}$ is the measurement error variance (“nugget”) and $\alpha$ is the ratio of the spatial variance (partial sill) to the total variance (sill).
Letting $\gamma^{{\mathrm{\scriptscriptstyle T}}}=\left[\beta^{{\mathrm{\scriptscriptstyle T}}},\Omega^{{\mathrm{\scriptscriptstyle T}}}\right]$ be $q\times(p+n)$, we assume $\{\gamma,\Sigma\}\sim\operatorname{MNIW}\left(\mu_{\gamma},V_{\gamma},\Psi_{0},\nu_{0}\right)$, where
$\operatorname{MNIW}\left(\gamma,\Sigma\mid\mu_{\gamma},V_{\gamma},\Psi_{0},\nu_{0}\right)=\operatorname{IW}(\Sigma\mid\Psi_{0},\nu_{0})\times\operatorname{MN}_{p,q}(\gamma\mid\mu_{\gamma}V_{\gamma},\Sigma),$ (2.5)
with $\mu_{\gamma}^{{\mathrm{\scriptscriptstyle T}}}=\left[m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0},0_{q\times n}\right]$ and $V_{\gamma}=$ blockdiag $\left\{M_{0},\rho_{\phi}(\mathcal{S},\mathcal{S})\right\}$. The mniw prior is conjugate with respect to the matrix-normal likelihood. Thus, for any fixed $\{\alpha,\phi\}$ and hyperparameters in the prior density, we obtain a MNIW posterior density for $\{\gamma,\Sigma\}$,
$p(\gamma,\Sigma\mid\mathscr{D})=\operatorname{MNIW}\left(\gamma,\Sigma\mid\mu_{\gamma}^{\star},V^{\star}_{\gamma},\Psi^{\star},\nu^{\star}\right),$ (2.6)
where $V^{\star}_{\gamma}=\left[\begin{array}[]{cc}\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}X+M_{0}^{-1}&\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}\\ \frac{\alpha}{1-\alpha}X&\rho_{\phi}^{-1}(\mathcal{S},\mathcal{S})+\frac{\alpha}{1-\alpha}\mathbb{I}_{n}\end{array}\right]^{-1}$ and $\mu_{\gamma}^{\star}=V^{\star}_{\gamma}\left[\begin{array}[]{c}\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}Y+m_{0}\\ \frac{\alpha}{1-\alpha}Y\end{array}\right]$, $\Psi^{\star}=\Psi_{0}+\frac{\alpha}{1-\alpha}Y^{{\mathrm{\scriptscriptstyle T}}}Y+m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0}m_{0}-\mu_{\gamma}^{*{\mathrm{\scriptscriptstyle T}}}V^{\star-1}_{\gamma}\mu_{\gamma}^{\star}$ and $\nu^{\star}=\nu_{0}+n$.
The framework in (2.4) is equivalent to (2.2) with $V=R_{\phi}+(\alpha^{-1}-1)\mathbb{I}_{n}$ with $R_{\phi}=[\rho(s_{i},s_{j};\phi)]$. We recover posterior samples of $\Omega$ by drawing a value of $\Omega$ from $p(\Omega\,|\,\mathscr{D},\beta,\Sigma,\mathscr{M}_{j})$ for every posterior draw of $\{\beta,\Sigma\}$. This renders itself seamlessly to the Bayesian transfer learning framework described in Section 2.1 provided that $V$, or $\{\alpha,\phi\}$, is fixed. For GeoAI, we seek to minimize human intervention. Rather than fixing them at one particular value, perhaps gleaned from a spatial variogram that requires human inspection, we use a set of $J$ candidate values $\{\alpha_{j},\phi_{j}\}$ specifying model $\mathscr{M}_{j}$ for $j=1,\ldots,J$. We now obtain analytical closed forms for $p(\beta,\Sigma\,|\,\mathscr{D},\mathscr{M}_{j})$ for each $j$, as described in Section 2.1, and use Bayesian predictive stacking to evaluate the stacked posterior distribution.
Turning to prediction, let $\mathcal{U}=\{u_{1},\dots,u_{n^{\prime}}\}$ be a finite set of locations where we seek to predict or impute the value of $Y$ based upon an observed ${n^{\prime}}\times p$ design matrix $X_{\mathcal{U}}$ associated with the locations in $\mathcal{U}$. The joint posterior predictive for $Y_{\mathcal{U}}$ and the unobserved latent process $\Omega_{\mathcal{U}}=[\omega(u_{i})^{{\mathrm{\scriptscriptstyle T}}}]$ for $i=1,\dots,n^{\prime}$, can be recast by integrating out $\{\gamma,\Sigma\}$ from the conditional posterior predictive distribution to yield
$\begin{split}p(Y_{\mathcal{U}},\Omega_{\mathcal{U}}\mid\mathscr{D})&=\int\mathrm{MN}_{{n^{\prime}},q}\big{(}Y_{\mathcal{U}}\mid X_{\mathcal{U}}\beta+\Omega_{\mathcal{U}},\left(\alpha^{-1}-1\right)\mathbb{I}_{{n^{\prime}}},\Sigma\big{)}\\ &\times\operatorname{MN}_{{n^{\prime}},q}\big{(}\Omega_{\mathcal{U}}\mid M_{\mathcal{U}}\Omega,V_{\Omega_{\mathcal{U}}},\Sigma\big{)}\times\operatorname{MNIW}\left(\gamma,\Sigma\mid\mu_{\gamma}^{\star},V^{\star}_{\gamma},\Psi^{\star},\nu^{\star}\right)d\gamma d\Sigma,\end{split}$ (2.7)
where $M_{\mathcal{U}}=\rho_{\phi}(\mathcal{U},\mathcal{S})\rho_{\phi}^{-1}(\mathcal{S},\mathcal{S})$ and $V_{\Omega_{\mathcal{U}}}=\rho_{\phi}(\mathcal{U},\mathcal{U})-\rho_{\phi}(\mathcal{U},\mathcal{S})\rho_{\phi}^{-1}(\mathcal{S},\mathcal{S})\rho_{\phi}(\mathcal{S},\mathcal{U})$. This is a matrix-variate Student’s t $\operatorname{T}_{2n^{\prime},q}(\nu^{\star},\mu^{\star},V^{\star},\Psi^{\star})$, with degrees of freedom $\nu^{\star}$, location matrix $\mu^{\star}=M\mu_{\gamma}^{\star}$, row-scale matrix $V^{\star}$, and column-scale matrix $\Psi^{\star}$, where $M=\begin{bmatrix}0&M_{\mathcal{U}}\\ X_{\mathcal{U}}&M_{\mathcal{U}}\end{bmatrix}$ and $V^{\star}=MV^{\star}_{\gamma}M^{\mathrm{\scriptscriptstyle T}}+V_{E}$ with $V_{E}=\begin{bmatrix}V_{\Omega_{\mathcal{U}}}&V_{\Omega_{\mathcal{U}}}\\ V_{\Omega_{\mathcal{U}}}&V_{\Omega_{\mathcal{U}}}+(\alpha^{-1}-1)\mathbb{I}_{n^{\prime}}\end{bmatrix}$. See Section LABEL:sec:matrixT for details. While the conditional posterior predictive distributions take the following forms $p(\Omega_{\mathcal{U}}\mid\mathscr{D},\gamma,\Sigma)=\operatorname{MN}_{{n^{\prime}},q}\left(\Omega_{\mathcal{U}}\mid M_{\mathcal{U}}\Omega,V_{\Omega_{\mathcal{U}}},\Sigma\right)$, and $p(Y_{\mathcal{U}}\mid\mathscr{D},\Omega_{\mathcal{U}},\gamma,\Sigma)=\operatorname{MN}_{{n^{\prime}},q}(Y_{\mathcal{U}}\mid X_{\mathcal{U}}\beta+\Omega_{\mathcal{U}},\left(\alpha^{-1}-1\right)\mathbb{I}_{{n^{\prime}}},\Sigma)$. Hence, we can proceed with posterior predictive inference by sampling from the closed-form joint predictive distribution or sampling from the conditional distributions. We draw one instance of $\Omega_{\mathcal{U}}\sim p(\Omega_{\mathcal{U}}\mid\mathscr{D},\gamma,\Sigma)$ for each posterior draw of $\{\gamma,\Sigma\}$ and then draw one value of $Y_{\mathcal{U}}\sim p(Y_{\mathcal{U}}\mid\mathscr{D},\Omega_{\mathcal{U}},\gamma,\Sigma)$ for each drawn $\left\{\Omega_{\mathcal{U}},\gamma,\Sigma\right\}$. The resulting samples are exactly drawn from the posterior predictive distribution $p(Y_{\mathcal{U}}\,|\,\mathscr{D})$.
This tractability is possible if the range decay $\phi$ and $\alpha$ are fixed. While data can inform about these parameters, they are inconsistently estimable and lead to poorer convergence (Zhang,, 2004). Finley et al., (2019) explored $K$-fold cross-validation, but inference is limited to only one set of values for the parameters. Instead, we pursue exact inference using (2.6) and (2.7) by stacking over different fixed values of $\{\alpha,\phi\}$ using bps of predictive densities as described in Section 2.2. This minimizes human intervention and enables automation.
For each subset of the data we compute the stacking weights $z_{k}=\{z_{k,j}\}_{j=1,\ldots,J}$ as
$\max_{z_{k}\in S_{1}^{J}}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\log\sum_{j=1}^{J}z_{k,j}p\left(Y_{k,i}\mid\mathscr{D}_{k,[-l]},\mathscr{M}_{j}\right),$ (2.8)
where $Y_{k,i}$ is the $i$-th of the $n_{k}$ rows of $Y_{k,[l]}\in\mathscr{D}_{k,[l]}$ (the $l$-th fold within the $k$-th dataset), and $\mathscr{D}_{k,[-l]}$ is the $k$-th dataset without the $l$-th fold, with $l=1,\dots,L$, and $L$ the number of folds for K-fold cross-validation estimates for the expected value of the score (see Section 2.2). The posterior predictive density, $p\left(Y_{k,i}\mid\mathscr{D}_{k,[-l]},\mathscr{M}_{j}\right)$, is available in closed form as a matrix t distribution, which makes the computation efficient. Leading Equation (2.8) to
$\max_{z_{k}\in S_{1}^{J}}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\log\sum_{j=1}^{J}z_{k,j}\operatorname{T}_{1,q}(Y_{k,i}\mid\nu^{\star}_{[-l]},\mu^{\star}_{i},V^{\star}_{i},\Psi^{\star}_{[-l]}),$ (2.9)
where $\nu^{\star}_{[-l]}=\nu_{0}+n_{k,[-l]}$, $n_{k,[-l]}$ is the cardinality of $\mathcal{S}_{k,[-l]}$ (which is the set of locations in $\mathscr{D}_{k,[-l]}$), and $\Psi^{\star}_{[-l]}=\Psi_{0}+(\alpha_{j}^{-1}-1)Y^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0}m_{0}-\mu^{{\mathrm{\scriptscriptstyle T}}\star}_{\gamma,[-l]}V^{-1\star}_{[-l]}\mu^{\star}_{\gamma,[-l]}$. While $V^{\star}_{i}=M_{y,i}V_{\gamma}^{\star}M_{y,i}^{{\mathrm{\scriptscriptstyle T}}}+V_{\Omega_{i}}+(\alpha_{j}^{-1}-1)$, $\mu^{\star}_{i}=M_{y,i}\mu^{\star}_{\gamma,[-l]}$, which are defined by computing the following auxiliary quantities: $V^{-1\star}_{\gamma,[-l]}=\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}X_{k,[-l]}+M_{0}^{-1}&\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}\\ \frac{\alpha_{j}}{1-\alpha_{j}}X_{k,[-l]}&\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})+\frac{\alpha_{j}}{1-\alpha_{j}}\mathbb{I}_{n_{k,[-l]}}\end{bmatrix}$, $\mu^{\star}_{\gamma,[-l]}=V^{\star}_{\gamma,[-l]}\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}\\ \frac{\alpha_{j}}{1-\alpha_{j}}Y_{k,[-l]}\end{bmatrix}$, $M_{y,i}=\begin{bmatrix}X_{k,i}&\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})\end{bmatrix}$, and $V_{\Omega_{i}}=\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,i})-\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})\rho_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,i})$. Note that $\nu^{\star}_{[-l]}$ is scalar, $\mu^{\star}_{i}$ is $(1\times q)$ row vector, $V^{\star}_{i}$ is a scalar, and $\Psi^{\star}_{[-l]}$ is $(q\times q)$ matrix. Further details, including derivations and implementation, are supplied in Section LABEL:sec:matrixT and Algorithm 1.
For each dataset $\mathscr{D}_{k}$, bps computes: (i) an estimate of the posterior predictive $\hat{p}(\cdot\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$ for $k=1,\dots,K$ and $j=1,\dots,J$; and (ii) a set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$. Once these weights are available, we apply bps a second time to obtain a weighted average of $\hat{p}(\cdot\mid\mathscr{D}_{k})$ over the $k$ subsets. This double bps of predictive densities seeks weights $w=\{w_{k}\}_{k=1,\ldots,K}$ such that $\hat{w}=\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}\hat{p}\left(Y_{k,i}\mid\mathscr{D}_{k,[-l]}\right)$ Once the stacking weights $\hat{w}=\{\hat{w}_{k}\}_{k=1,\dots,K}$ are obtained, the double bps estimation of any posterior or posterior predictive distributions of interest is obtained from
$\hat{p}(\cdot\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)\;.$ (2.10)
Only posterior predictive distributions are considered to acquire the two sets of stacking weights. Given the weights obtained from double stacking, the stacked posterior distribution is a mixture of finite mixtures. This makes sampling from (2.10) straightforward.
First, the set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$ obtained using bps within the subset of the data $\mathscr{D}_{k}$, is primarily used to approximate the subset posterior distribution $\hat{p}(\Theta\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}p\left(\Theta\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$, for each subset $k=1,\dots,K$. By considering the second set of stacking weights $\hat{w}=\{\hat{w_{k}}\}_{k=1,\dots,K}$, the stacked full posterior distribution is $\hat{p}(\Theta\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\hat{p}(\Theta\mid\mathscr{D}_{k})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p(\Theta\mid\mathscr{D}_{k},\mathscr{M}_{j})$, where $\Theta=\{\gamma,\Sigma\}$. This is a substantial simplification over meta-kriging (Guhaniyogi and Banerjee,, 2018, 2019) that does not require any empirical approximation of posterior or predictive distributions. Again, the predictive random variable $Y_{\mathcal{U}}$ is recovered from (2.10) as
$\hat{p}(Y_{\mathcal{U}}\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(Y_{\mathcal{U}}\mid\mathscr{D}_{k},\mathscr{M}_{j}\right).$ (2.11)
Inferential interest resides with the posterior predictive surface for $\Omega$. Thus, we estimate $\Omega_{\mathcal{U}}$ using $\hat{p}(\Omega_{\mathcal{U}}\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\hat{p}(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k})$, where $\hat{p}(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$.
Algorithm 1 Computing stacking weights within subsets using bps
Input: $Y$ ($n\times q$ matrix of outcomes), $X$ ($n\times p$ design matrix), $\mathcal{S}$ (coordinates of $n$ locations); $\{m_{0},M_{0},\Psi_{0},\nu_{0}\}$: Prior parameters; $G_{\alpha}\times G_{\phi}$: Grids of $\{\alpha,\phi\}$; $n$ (no. of locations), $q$ (no. of outcomes), $p$ (no. of predictors); $K$ (no. of subsets), $J$ (no. of models), $L$ (no. of folds).
Output: $\hat{z}=\{\hat{z}_{k}=\{\hat{z}_{k,j}\}:k=1,\dots,K,j=1,\dots,J\}$: Stacking weights within subsets; $\{pd_{k,j,i}:k=1,\dots,K,j=1,\dots,J,i=1,\dots,n\}$: point-wise predictive density of $Y$; $G_{all}$: Grid of dimension $J$, spanned by $G_{\alpha},G_{\phi}$
1:Partition $Y,X,\mathcal{S}$ into $\mathscr{D}_{k}=\{Y_{k},X_{k},\mathcal{S}_{k}\}$, $k=1,\dots,K$
2:Store $n_{k}$, as cardinality of $\mathcal{S}_{k}$; Compute $G_{all}$ by expanding $G_{\alpha},G_{\phi}$
3:for $k=1,\dots,K$ do Parallel
4:     for $j=1,\dots,J$ do
5:         Extract $\{\alpha_{j},\phi_{j}\}$ from $j$-th row of $G_{all}$
6:         Form $L$ folds: $\mathscr{D}_{k,[l]}=\{Y_{k,[l]},X_{k,[l]},\mathcal{S}_{k,[l]}\}$ and $\mathscr{D}_{k,[-l]}=\{Y_{k,[-l]},X_{k,[-l]},\mathcal{S}_{k,[-l]}\}$
7:         Store $n_{k,[-l]}$, as cardinality of $\mathcal{S}_{k,[-l]}$
8:         for $l=1,\dots,L$ do
9:              Compute $R_{\phi_{j}}([-l])=\rho_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})$, $R^{-1}_{\phi_{j}}([-l])$ and $M_{0}$ for $M_{0}^{-1}$
10:              Construct $V^{-1\star}_{\gamma,[-l]}=$$\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}X_{k,[-l]}+M_{0}^{-1}&\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}\\ \frac{\alpha_{j}}{1-\alpha_{j}}X_{k,[-l]}&R^{-1}_{\phi_{j}}([-l])+\frac{\alpha_{j}}{1-\alpha_{j}}\mathbb{I}_{n_{k,[-l]}}\end{bmatrix}$
11:              Solve for $\mu^{\star}_{\gamma,[-l]}$: $V^{-1\star}_{\gamma,[-l]}\mu^{\star}_{\gamma,[-l]}=\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}\\ \frac{\alpha_{j}}{1-\alpha_{j}}Y_{k,[-l]}\end{bmatrix}$
12:              Calculate $\Psi^{\star}_{[-l]}=\Psi_{0}+(\alpha_{j}^{-1}-1)Y^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0}m_{0}-\mu^{{\mathrm{\scriptscriptstyle T}}\star}_{\gamma,[-l]}V^{-1\star}_{[-l]}\mu^{\star}_{\gamma,[-l]}$
13:              Calculate $\nu^{\star}_{[-l]}=\nu_{0}+n_{k,[-l]}$
14:              for $i\in[l]$ do
15:                  Compute $R_{\phi_{j}}(i)=\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,i})$ and $R_{\phi_{j}}(i,[-l])=\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})$
16:                  Calculate $M_{i}=R_{\phi_{j}}(i,[-l])R^{-1}_{\phi_{j}}([-l])$ and form $M_{y,i}=\begin{bmatrix}X_{k,i}&M_{i}\end{bmatrix}$
17:                  Calculate $\mu^{\star}_{i}=M_{y,i}\mu^{\star}_{\gamma,[-l]}$ and $V_{\Omega_{i}}=R_{\phi_{j}}(i)-M_{i}R_{\phi_{j}}([-l],i)$
18:                  Construct $V_{e,i}=V_{\Omega_{i}}+(\alpha_{j}^{-1}-1)$ and $V^{\star}_{i}=M_{y,i}V^{\star}_{\gamma,[-l]}M_{y,i}^{{\mathrm{\scriptscriptstyle T}}}+V_{e,i}$
19:                  Compute $pd_{k,j,i}=\operatorname{T}_{1,q}(Y_{k,i}\mid\nu^{\star}_{[-l]},\mu^{\star}_{i},V^{\star}_{i},\Psi^{\star}_{[-l]})$.
20:              end for
21:         end for
22:     end for
23:     Solve: $\max_{z_{k}\in S_{1}^{J}}\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}\log\sum_{j=1}^{J}z_{k,j}pd_{k,j,i}$ such that $z_{k}\in[0,1]^{J}:\sum_{j=1}^{J}z_{k,j}=1$
24:end Parallel for
25:return $\{\hat{z},\{pd_{k,j,i}\},G_{all}\}$
3 Computer programs and resources
Our subsequent implementations, including all simulations experiments and data analysis, are executed in native R and c++ deploying the spBPS package. In particular, all programs used to obtain the ensuing analysis are publicly accessible from the GitHub repository lucapresicce/Bayesian-Transfer-Learning-for-GeoAI that links the Rcpp-based spBPS package including the main function executing bps for univariate and multivariate geostatistical models. The reported results are executed on a standard laptop, which runs an Intel Core I7-$8750$H CPU with $5$ available cores for parallel computation and $16$ GB of RAM.
We conduct experiments comparing double bps with spatial meta-kriging (smk) using a meta-posterior (Minsker et al.,, 2017) defined as the geometric median of individual posterior distributions based on subsets of the data (as implemented in Guhaniyogi and Banerjee,, 2018, 2019). We execute meta-kriging using the spMvLM function within the spBayes R package that fits a linear model of coregionalization for multivariate spatial data using MCMC (Finley et al.,, 2015). For parallel implementations of bps and smk, we employ R packages doParallel, and foreach (Corporation and Weston,, 2022; Microsoft and Weston,, 2022). We map the interpolated spatial surfaces using the MBA R package (Finley et al.,, 2022), while sampling from the matrix-variate normal and t distributions are achieved using the mvnfast R package (Fasiolo,, 2014). Section LABEL:sec:compdetails in the Supplement provides further details including memory management considerations and efficient solutions to spatial “BIG” data analysis, and sensitivity on the number of data shards $K$.
Algorithm 2 Calculating stacking weights between subsets using bps
Input: $\hat{z}=\{\hat{z}_{k}=\{\hat{z}_{k,j}\}:k\in\{1,\dots,K\},j\in\{1,\dots,J\}\}$: Stacking weights within subsets; $\{pd_{k,j,i}:k=1,\dots,K,j=1,\dots,J,i=1,\dots,n\}$: point-wise predictive density of $Y$; $n,q,p$: Number of rows, number of outcomes, and number of predictors; $K,\{n_{k}:k\in\{1,\dots,K\}\},J$: Number of subsets, dimension of each subset, and number of competitive models in each subset.
Output: $\hat{w}=\{\hat{w}_{k}:k=1,\dots,K\}$: Stacking weights between subsets.
1:Construct $pd=[pd_{1}^{{\mathrm{\scriptscriptstyle T}}}:\cdots:pd_{K}^{{\mathrm{\scriptscriptstyle T}}}]^{{\mathrm{\scriptscriptstyle T}}}$ of dimension $(n\times J)$
where $pd_{k}=\begin{bmatrix}pd_{k,1,1}&\dots&pd_{k,J,1}\\ \vdots&pd_{k,j,i}&\vdots\\ pd_{k,1,n_{k}}&\dots&pd_{k,J,n_{k}}\\ \end{bmatrix}$ of dimension $(n_{k}\times J)$
2:for $k=1,\dots,K$ do
3:     Compute $epd_{k}=pd\;\hat{z}_{k}$ of dimension $(n\times 1)$
4:end for
5:Solve convex optimization problem:
$\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}epd_{k,i}=\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}\sum_{j=1}^{J}\hat{z}_{k,j}pd_{k,j,i}$
where $pd_{k,j,i}=p\left(Y_{k,i}\mid\mathscr{D}_{k,[-l]},\mathscr{M}_{j}\right)=\operatorname{T}_{1,q}(Y_{k,i}\mid\nu^{\star}_{[-l]},\mu^{\star}_{i},V^{\star}_{i},\Psi^{\star}_{[-l]})$
for $\forall i\notin[-l],l\in\{1,\dots,L\}$ and $S_{1}^{K}=\{w\in[0,1]^{K}:\sum_{k=1}^{K}w_{k}=1\}$
6:return $\hat{w}=\{\hat{w}_{k}:k\in\{1,\dots,K\}\}$
4 Simulation experiments
We conduct multiple simulation experiments to evaluate inferential performance of double bps while underscoring comparisons with smk. We present a selection of simulation results for multivariate models here and refer the reader to Section LABEL:sec:sim_suppl for further experiments.
4.1 Simulation - BPS vs SMK
In the first setting, we consider two synthetic datasets with common structures but distinct sizes using the model (2.4). Both datasets consist of $p=2$ predictors and $q=2$ response variables, but different numbers of spatial locations, $n=5,000$ and $n=10,000$, respectively. We generate spatial coordinates from a uniform distribution on $[0,1]^{2}$ domain and fix them. We build the $n\times n$ spatial correlation matrix $V$ over these coordinates using an exponential spatial correlation function with $\phi=4$ and specify $\Sigma=\mathbb{I}_{q}$. From these specifications we generate the $n\times q$ matrix $Y$ from the first equation of (2.4) with fixed $p\times q$ matrix $\beta=\begin{bmatrix}-0.75&1.85\\ 0.90&-1.10\end{bmatrix}$, a fixed $n\times p$ matrix $X$ with a first column of ones, representing the intercept, and $p-1$ columns of values randomly simulated from a uniform distribution on $[0,1]$ (emulating standardized predictors), and the proportion of spatial variability $\alpha=0.8$.
We analyze the simulated datasets using double bps and smk in two settings: (i) $K=10$ and $n=5,000$; (ii) $K=10$ and $n=10,000$; (iii) $K=5$ and $n=5,000$; and (iv) $K=20$ and $n=10,000$. These settings produce subsets of size $n/K\in\{500,1000\}$. We implement bps using $J=9$ candidate models $\mathscr{M}_{j},j=1,\dots,J$, where each model is specified by a set of candidate values for the hyperparameters ${\alpha_{j},\phi_{j}}$ in (2.4). These hyperparameters represent the proportion of spatial variability and the parameter(s) of the spatial correlation function, respectively. The set of candidate models is constructed as the set of all possible combinations of values for these hyperparameters. In the subsequent experiments, the grid of models was built using $\alpha\in\{0.70,0.80,0.90\}$ and $\phi\in\{3,4,5\}$. These values resemble an effective spatial range of $\{0.99,0.75,0.60\}$ units, corresponding to $70\%,53\%,42\%$ of the maximum inter-site distance inside the unit square, beyond which the spatial correlation drops below $0.05$. Equation (2.5) follows choices in Zhang et al., (2021). Specifically, we set $m_{0}=0_{p\times q}$, $M_{0}=10\mathbb{I}_{p}$, $\Psi_{0}=\mathbb{I}_{q}$, and $\nu_{0}=3$ in the $\operatorname{MNIW}$ joint prior for $\{\gamma,\Sigma\}$ in 2.5. For data analysis, we use an exponential spatial correlation function for $\rho_{\phi}(\cdot,\cdot)$, which is completely defined conditionally on $\mathscr{M}_{j}$, since specifies a value for $\phi$. Again, in the conjugate framework, we draw $R=250$ posterior samples used for inference.
Model $n=5000,K=10$ $n=5000,K=5$ $n=10000,K=20$ $n=10000,K=10$
smk 51.41 237.23 103.36 446.01
bps 1.38 7.22 2.24 10.58
Relative ratio 37.25 32.85 46.14 42.15
Table 1: Computational aspects for the different settings, with time in minutes.
We also apply smk to the two simulated datasets with the same combinations of $n$ and $K$ as bps. Unlike bps, where we stack analytically tractable posteriors over a range of fixed values of spatial covariance kernel parameters, the smk implementation attempts full Bayesian inference using prior distributions on spatial covariance kernel parameters. We fit the linear model of coregionalization described in Finley et al., (2015) for each subset of the multivariate spatial data using MCMC. The posterior samples from the $K$ subsets are combined using Weiszfeld’s iterative algorithm (Minsker et al.,, 2017) to produce an estimate of the geometric median of the posterior distributions.
Table 1, which compares the computational speed of double bps with smk reveals massive computational gains accrued from bps. The computational advantage evinced from the relative ratio become more pronounced as the size of the data becomes larger despite the larger subsets. This is explained by the fact that fitting the Gaussian process regression dominates the computation relative to the assimilation of inference from the subsets. If the number of locations explode, then the geometric median of posteriors required by smk is computationally unfeasible. While smk offers Bayesian estimates using MCMC for each subset, double bps avoids MCMC and, hence, issues of convergence.
Importantly, as we see in Figure 2, which depicts estimated response surfaces using bps and smk corresponding to $n=5000$ and $K=10$, inference from smk and bps are practically indistinguishable. The root mean squared prediction errors, rmspe, reported in Figure 2 denotes the average squared differences between the generated and estimated values of the response and reveals minor discrepancies between bps and smk. Section LABEL:sec:sim_suppl_M presents results for the other configurations of $n$ and $K$, each of which reveals that bps offers practically indistinguishable spatial interpolation from smk at a fraction of the computational cost.
Figure 2: from left to right: comparison between the true generated response surfaces, the surfaces predicted from bps and smk (posterior mean), with rmspe. For $n=5000$, $K=10$.
Figure 3 reports 95% posterior predictive intervals for the response. Again, the empirical coverage is impressive. While we see slightly wider intervals from BPS, this is less pronounced than the underestimation of variability seen with smk. Moreover, Figure 3 reveals superior map estimates for the bps. Finally, Figure 4 presents the recovery of parameter estimates. As seen in predictive inference, the posterior credible intervals for parameters also deliver practically indistinguishable inference for the two modeling frameworks. In particular, both methods recover the true values for $\beta$ and $\Sigma$, while BPS reconstructs a better point estimate for range parameters $\phi$ using $\sum_{k=1}^{K}\hat{w}_{k}\sum_{j=1}^{J}\hat{z}_{j}\phi_{j}$. Similar conclusions are drawn from other values of $n$ and $K$ explored in Section LABEL:sec:sim_suppl for univariate responses.
Figure 3: from top to bottom: comparison between posterior predictive intervals for the predicted response from bps and smk, with empirical coverage. For $n=5000$, $K=10$.
Figure 4: from left to right: comparison between posterior credible intervals for the parameters recovered from bps and smk. For $n=5000$, $K=10$.
4.2 Simulation - Multivariate Transfer learning
We now consider inferential performance for different misspecified models. As in the earlier example, we generate the $n\times q$ outcome $Y$ using  (2.4) with $n=5,000$, $q=3$ and $p=2$, $X$ includes an intercept and $p-1$ columns generated from a standard uniform distribution over $[0,1]$, $\beta=\begin{bmatrix}-0.75&1.05&-0.35\\ 2.20&-1.10&0.45\end{bmatrix}$ and $\Sigma=\begin{bmatrix}2.00&0.80&0.20\\ 0.80&2.00&-0.45\\ 0.20&-0.45&2.00\end{bmatrix}$. The $n\times n$ spatial correlation matrix $V$ is specified using an exponential correlation matrix with a fixed range parameter at $\phi=4$ and $\alpha=0.8$ to be the strength of spatial dependence.
We analyze the generated data using three different models that assign prior distributions to $\{\beta,\Sigma\}$ with $m_{0}=0_{p\times q}$, $M_{0}=10\mathbb{I}_{p}$, $\Psi_{0}=\mathbb{I}_{q}$, and $\nu_{0}=3$, but different values of $\alpha$ and $\phi$ representing (i) well-specified setting with the data generating values $\{\alpha=0.8,\phi=4\}$; (ii) moderately misspecified setting with $\{\alpha=0.45,\phi=6.63\}$; and (iii) a highly misspecified setting with $\{\alpha=0.25,\phi=50\}$. For each setting, we partition the original data into $K=10$ subsets, corresponding to subsets of size $500$. We consider $J=9$ competitive models for bps and consider $\alpha\in\{0.75,0.80,0.85\}$ and $\phi\in\{2,4,6\}$ that yields effective spatial ranges in the percentage of maximum point inter-distance of $105\%,53\%,35\%$ respectively.
Here, we display results for the bps, while graphical results for other settings are available in Section LABEL:sec:sim_suppl_M. Figure 5 presents a visual comparison between the generated true response and the posterior mean of spatial predictions delivered by bps. We present the rmspe along with a set of credible intervals for the parameters relative to the true values. The inferential performance evaluated using the rmspe as well as the spatially interpolated maps corroborates the effectiveness of bps. Figure 6 reveals that more than half of the model unknowns are included in the 95% posterior credible interval. Unsurprisingly, inferential performance is weakest for the spatial variance, which is not identifiable from the realized data (Zhang,, 2004). Further results on simulation experiments are available in Section LABEL:sec:sim_suppl_M.
Figure 5: from top to bottom: comparison between the true generated spatial surfaces and the posterior mean surfaces of predictions from bps, with rmspe for each outcome.
Figure 6: Posterior credible intervals for parameters recovered by bps, with true parameters in yellow.
Table 2 presents predictive performance in terms of rmspe, along with the hyperparameter values corresponding to the different experimental settings. The first three rows correspond to fixed values of $\{\alpha,\phi\}$ in the Bayesian transfer learning approach of Section 2.1. The last row corresponds to bps when $\alpha$ and $\phi$ are weighted averages using the stacking weights as described in the last paragraph of Section 4.1. While bps outperforms transfer learning in terms of rmspe, some remarks on transfer learning are merited. The transfer learning approach offers slightly better predictive performance than the well-specified model with misspecified data than with well-specified data. This, while somewhat unexpected, is attributed to the estimated response surfaces being practically indistinguishable. Furthermore, transfer learning in even the highly misspecified setting performs competitively. In conclusion, these findings seem to be consistent with theoretical insights offered by Stein, (1988) and Stein and Handcock, (1989), who theoretically established the ability of Gaussian processes to deliver good predictive performance even for misspecified covariance functions within the in-fill asymptotic paradigm.
Model $\alpha$ $\phi$ Average rmspe
well specified 0.8 4 0.942
misspecified 0.45 6.63 0.963
highly misspecified 0.25 50 1.236
double bps 0.75 4.11 0.900
Table 2: Average rmspe and specification for each model scenario.
5 Application
Global warming and its critical consequences (see Fisher,, 1958; Nicholls,, 1989; Friehe et al.,, 1991; O’Carroll et al.,, 2019) are increasingly being investigated using machine learning and AI tools that assist in developing predictive global models and formulating data-driven policies. trying to produce faster solutions. The key aspects of global warming are related to global temperature, atmospheric and land surface compositions. We analyze data extracted from NASA’s Moderate Resolution Imaging Spectroradiometer (modis) platform.
Specifically, we consider as response variables two vegetation indices present in vi data: the normalized difference vegetation index (ndvi); and the red reflectance (rr). Both measure the activity of vegetation on land surfaces and are commonly investigated to help understand biophysical and structural properties of vegetation at global scales (Walther et al.,, 2018; Jiang et al.,, 2020; Ai et al.,, 2020). The predictors comprise an intercept and the solar zenith angle for that location ($p=2$). The latter depicts the strength of solar irradiation. All the data were collected over a 16-day window in May 2024 and then averaged.
[…]
The multivariate model developed in Section 2.3 with ndvi and rr comprising the $q=2$ columns of $Y$. Figures 7 and 8 illustrate maps corresponding to ndvi and rr, respectively, using $K=2,000$ subsets. The top left panel in each figure presents the spatially interpolated map of the training data for the respective responses. The interpolated maps for the test data are displayed in the top right panel of the respective figures. These maps reveal pronounced spatial variation, where the darker shades of green in ndvi represent higher values of detected biomass, while lighter shades of brown represent low biomass. Conversely for rr, warmer colors in the red spectrum represent higher reflectance.
Figure 7: Left to right: Maps for training data (top left), test data (top right) and predicted surface (bottom right) for ndvi. Empirical coverage for held-out values are in the bottom left. Results correspond to $K=2,000$.
Figure 8: Left to right: Maps for training data (top left), test data (top right) and predicted surface (bottom right) for rr. Empirical coverage for held-out values of outcomes (bottom left). Results correspond to $K=2,000$.
Figures 7 and 8 show results from double bps. Maps of interpolated posterior means and predicted responses are displayed in the bottom right of the respective figures. The bps interpolated maps are almost indistinguishable from the true surface (top right). This suggests that the automated bps effectively, perhaps even strikingly, recovers the spatial patterns in spite of the modeling simplifications over more elaborate statistical models (Banerjee,, 2020; Zhang et al.,, 2021). Moreover, the double bps behaves well in terms of predictive performance and uncertainty quantification for both values of $K$ considered. It achieves an rmspe of $(0.206,0.159)$ when $K=2,000$, and $(0.241,0.197)$ with $K=4,000$ subsets for ndvi and rr respectively. This is significantly lower (by almost 100%) than the standard Bayesian conjugate multivariate linear model, which produces rmspe of $(0.474,0.439)$ for the two vegetation indexes, respectively. The predictive $95\%$ credible intervals when $K=2,000$ for both responses are shown in the bottom left of Figures 7 and 8, respectively.
[…]
Conjugate Linear model bps ($K=4,000$) bps ($K=2,000$)
$\beta_{0,\textsc{ndvi}}$ 34.495 ( 34.392, 34.601) 1.364 (-1.039, 3.907) 1.767 (-0.422, 3.917)
$\beta_{1,\textsc{ndvi}}$2.708 ( -2.719, -2.697) 0.744 ( 0.478, 0.998) 0.707 ( 0.485, 0.939)
$\beta_{0,\textsc{rr}}$16.924 (-17.014, -16.837) -0.418 (-2.470, 1.710) -0.562 (-2.187, 0.962)
$\beta_{1,\textsc{rr}}$ 2.664 ( 2.655, 2.674) 0.945 ( 0.726, 1.158) 0.959 ( 0.795, 1.139)
$\Sigma_{\textsc{ndvi}}$ 0.221 ( 0.220, 0.221) 0.182 ( 0.130, 0.257) 0.119 ( 0.099, 0.146)
$\Sigma_{\textsc{ndvi},\textsc{rr}}$0.167 ( -0.168, -0.166) -0.114 (-0.173, -0.079) -0.076 (-0.094, -0.061)
$\Sigma_{\textsc{rr}}$ 0.155 ( 0.154, 0.155) 0.115 ( 0.082, 0.170) 0.077 ( 0.062, 0.092)
$\varrho_{\textsc{ndvi},\textsc{rr}}$0.921 ( -0.928, -0.915) -0.816 (-0.874, -0.706) -0.844 (-0.880, -0.795)
rmspe [0.474, 0.404, 0.439] [0.241, 0.197, 0.219] [0.206, 0.159, 0.183]
Time 12 21 81
Table 3: Vegetation Index data analysis parameter estimates, rmspe, and computing time in minutes for candidate models. Parameter posterior summary 50 (2.5, 97.5) percentiles. Root mean square prediction error(s) presentation [ndvi, red reflectance, average].
Recalling that all the analyses are produced on a personal laptop (with just 5 physical cores) with minimal human intervention, the total run time of only 21 minutes with $K=4,000$ and 81 minutes with $K=2,000$ are impressive confirm the quadratic dependence of the partition size discussed in Section LABEL:sec:data1. Moreover, the strong dependence on the number of $J$ competitive models assessed is worth noting. Here $J=4$ while in Section LABEL:sec:data1 $J=3$. This suggests a marginal computational burden for each competitive model.
6 Discussion
This manuscript lays out a statistical modeling component for an artificially intelligent geospatial system (GeoAI) tailored to handle massive spatial datasets with minimal human intervention. The contribution harnesses analytically accessible multivariate statistical distributions in conjunction with Bayesian predictive stacking to deliver rapid inference on multivariate processes and circumventing iterative algorithms that require, often extensive, human tuning or control. Our proposed GeoSpatial AI system is capable of handling truly massive amounts of spatial data that preclude accessibility even on high performance CPU architectures and we develop Bayesian transfer learning using double Bayesian predictive stacking.
A few additional remarks are warranted. The development here has been elucidated with a hierarchical matrix-variate spatial process framework. While modeling simplifications have been introduced to minimize human intervention, we wish to emphasize that the bps seamlessly applies to more versatile but analytically intractable models. First, for spatially misaligned multivariate data where not all the variables may have been observed over the same set of locations, we encounter missing entries in $Y$ but are still able to deliver analytically tractable closed-form distribution theory (Zhang and Banerjee,, 2022). Second, should users wish to explore more complex models with unknown $\{\alpha,\phi\}$ or work with alternate multivariate models, then we can implement full Bayesian inference (now using MCMC, Variational Bayes or Laplace approximations) on each subset and invoke bps to pool across the subsets of the data. Time for analyzing each subset increase due to model complexity, but scaling inference to massive datasets is still enabled by bps—and that too with considerable efficiency over alternate divide-and-conquer approaches such as the geometric median, or gm, of posteriors (used in meta-kriging developed by Guhaniyogi and Banerjee,, 2018).
A preliminary examination of Section 2.3 reveals that posterior inference is conceptualized as a “mixture of mixtures”. Following Yao et al., (2018), we prefer stacking to fitting a mixture model because the former is numerically robust and requires almost no human tuning. Moreover, the structure of mixture-of-mixtures conceptually resembles the “Mixture of Experts” (MoE) that are adopted by AI platforms such as GPT-4 and Mistral. Disseminating our proposed product with our accompanying software, which is currently being migrated to R, is expected to significantly boost Geospatial AI systems. Future directions will also explore perceived potential of bps as a feeder for emerging amortized inference methods (Ganguly et al.,, 2023; Zammit-Mangion et al.,, 2024; Sainsbury-Dale et al.,, 2024) to achieve Bayesian inference. Rapid delivery of posterior estimates of the entire spatial process from bps will amount to more training data for amortized neural learners that can result in accelerated tuning for subsequent Bayesian inference. We do not see our proposed approach as a competitor to, but rather as supplementary to amortized neural inference. Such developments will be pursued as future research. We also seek to expand and fully investigate automated double bps using Markovian graphical structures across data subsets to further expedite and improve Geospatial AI systems.
Supplementary Materials
The online supplement includes additional derivations and details on matrix-variate distributions (Section LABEL:sec:distributiontheory), computational details and algorithms (Section LABEL:sec:compdetails), additional simulation experiment results (Section LABEL:sec:sim_suppl), details on accelerated bps for univariate spatial data (Section LABEL:sec:AccelerUnivarMod) and analysis of an additional dataset on sea surface temperature (Section LABEL:sec:data1). Section LABEL:sec:eda outlines automated computation of model hyperparameters without human intervention. Computer programs to reproduce all experiments and data analysis are publicly accessible from the GitHub repository lucapresicce/Bayesian-Transfer-Learning-for-GeoAI

### v2

@@
 This article devises transfer learning frameworks for deployment in artificially intelligent systems, where a massive data set is split into smaller data sets that stream into the analytical framework to propagate learning and assimilate inference for the entire data set.
-Specifically, we introduce Bayesian predictive stacking for multivariate spatial data and demonstrate its effectiveness in rapidly analyzing massive data sets.
-Furthermore, we make inference feasible in a reasonable amount of time, and without excessively demanding hardware settings.
-We illustrate the effectiveness of this approach in extensive simulation experiments and subsequently analyze massive data sets in climate science on sea surface temperatures and on vegetation index.
+Specifically, we introduce Bayesian predictive stacking for multivariate spatial data and demonstrate rapid and automated analysis of massive data sets.
+Furthermore, inference is delivered without human intervention without excessively demanding hardware settings.
+We illustrate the effectiveness of our approach through extensive simulation experiments and in producing inference from massive datasets on sea surface temperatures and vegetation index that are indistinguishable from traditional (and more expensive) statistical approaches.
 Keywords— Bayesian predictive stacking; Climate science; Gaussian process; Geospatial systems; Matrix-variate spatial models; Transfer learning.
@@
 We first obtain stacked inference for each of the parameters within each subset and, then, assimilate inference across subsets by a second stacking algorithm.
-A flowchart is supplied in Figure 1 comprising three steps: (i) partition the data into smaller subsets; (ii) analyze each subset by stacking conjugate posteriors in closed form; and (iii) stack across data sets to deliver inference for the entire data.
+A flowchart is supplied in Figure 1 comprising three steps: (i) partition data into smaller subsets; (ii) analyze each subset by stacking conjugate posteriors in closed form; and (iii) stack across data sets to deliver inference for the entire data.
 […]
@@
 Section 5 analyzes a very large climate data set with observed locations in the millions at global scales.
-Finally, Section 6 concludes the paper with some discussion.
+Finally, Section 6 concludes with some discussion.
+References to an accompanying Supplement are provided throughout the article and after Section 6.
 2 Bayesian transfer learning
@@
 Once these weights are available, we apply bps a second time to obtain a weighted average of $\hat{p}(\cdot\mid\mathscr{D}_{k})$ over the $k$ subsets.
-This double bps of predictive densities seeks weights $w=\{w_{k}\}_{k=1,\ldots,K}$ such that $\hat{w}=\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}\hat{p}\left(Y_{k,i}\mid\mathscr{D}_{k,[-l]}\right)$ Once the stacking weights $\hat{w}=\{\hat{w}_{k}\}_{k=1,\dots,K}$ are obtained, the double bps estimation of any posterior or posterior predictive distributions of interest is obtained from
+This double bps of predictive densities seeks weights $w=\{w_{k}\}_{k=1,\ldots,K}$ such that $\hat{w}=\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}\hat{p}\left(Y_{i}\mid\mathscr{D}_{k}\right)$ (see Section LABEL:sec:BPSobjfun).
+Once the stacking weights $\hat{w}=\{\hat{w}_{k}\}_{k=1,\dots,K}$ are obtained, the double bps estimation of any posterior or posterior predictive distributions of interest is obtained from
 $\hat{p}(\cdot\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)\;.$ (2.10)
@@
 6 Discussion
-This manuscript lays out a statistical modeling component for an artificially intelligent geospatial system (GeoAI) tailored to handle massive spatial datasets with minimal human intervention.
+This manuscript devises a statistical modeling component for an artificially intelligent geospatial system (GeoAI) tailored to analyze massive datasets with minimal human intervention.
 The contribution harnesses analytically accessible multivariate statistical distributions in conjunction with Bayesian predictive stacking to deliver rapid inference on multivariate processes and circumventing iterative algorithms that require, often extensive, human tuning or control.

### v3

@@
-supplement_jasa
+Bayesian Transfer Learning for Artificially Intelligent Geospatial Systems: A Predictive Stacking Approach
 Abstract
-Building artificially intelligent geospatial systems require rapid delivery of spatial data analysis at massive scales with minimal human intervention.
-Depending upon their intended use, data analysis may also entail model assessment and uncertainty quantification.
+Building artificially intelligent geospatial systems requires rapid delivery of spatial data analysis on massive scales with minimal human intervention.
+Depending upon their intended use, data analysis can also involve model assessment and uncertainty quantification.
 This article devises transfer learning frameworks for deployment in artificially intelligent systems, where a massive data set is split into smaller data sets that stream into the analytical framework to propagate learning and assimilate inference for the entire data set.
@@
 Furthermore, inference is delivered without human intervention without excessively demanding hardware settings.
-We illustrate the effectiveness of our approach through extensive simulation experiments and in producing inference from massive datasets on sea surface temperatures and vegetation index that are indistinguishable from traditional (and more expensive) statistical approaches.
+We illustrate the effectiveness of our approach through extensive simulation experiments and in producing inference from massive dataset on vegetation index that are indistinguishable from traditional (and more expensive) statistical approaches.
 Keywords— Bayesian predictive stacking; Climate science; Gaussian process; Geospatial systems; Matrix-variate spatial models; Transfer learning.
 1 Introduction
-Geospatial artificial intelligence (GeoAI) is a rapidly evolving discipline at the interface of machine learning and spatial data science that attempts to harness the analytical capabilities of Artificial Intelligence (AI) to analyze massive amounts of geographic data and conduct data driven scientific discoveries in diverse domains within the environmental and physical sciences.
+Geospatial artificial intelligence (GeoAI) is a rapidly evolving discipline at the interface of machine learning and spatial data science that attempts to harness the analytical capabilities of Artificial Intelligence (AI) to analyze massive amounts of geographic data for data driven scientific discoveries in the environmental and physical sciences.
 The area, while still fledgling, presents opportunities for statisticians to devise essential data analytic tools that should comprise an artificially intelligent “geospatial” system.
-Constructing comprehensive AI-based geospatial systems involve expertise in computer science relating to efficiently designed Geographic Information Systems (GIS) databases, conventional machine learning tools based upon algorithms for spatial interpolation, and statistical or probabilistic learning tools to offer scientific inference.
 This manuscript concerns itself primarily with statistical learning tools for artificially intelligent GeoAI systems.
-What, if at all, should be the role of formal statistical inference in GeoAI?
-Formal inference for spatial and spatial-temporal random fields enjoy a venerable intellectual presence spanning several decades of theoretical developments within classical and Bayesian paradigms (see, e.g., Cressie,, 1993; Stein,, 1999; Gelfand et al.,, 2010; Cressie and Wikle,, 2011; Banerjee et al.,, 2015).
-With increasing complexity in spatially oriented datasets, statisticians have built richly structured hierarchical models in their pursuit of “full” probabilistic inference.
-The term “full” loosely refers to approaches that attempt to estimate all unknowns in the model including all parameters (irrespective of how well they are identified by the data), random effects or partial realizations of the process, and predictive random variables for probabilistic interpolation at arbitrary locations (and time points in spatial-temporal data).
-Statisticians would, we believe, broadly agree that GeoAI offers sufficient intellectual space to accommodate rigorous probabilistic learning as a crucial constituent in artificially intelligent data analysis systems.
+What should be the role of statistical inference in GeoAI?
+Formal inference for spatial random fields enjoy a venerable intellectual presence spanning several decades of theoretical developments within classical and Bayesian paradigms (see, e.g., Cressie, , 1993; Stein, , 1999; Gelfand et al., , 2010; Cressie and Wikle, , 2011; Banerjee et al., , 2015).
+Statisticians have built richly structured hierarchical models in pursuit of “full” probabilistic inference for richly structured spatial data.
+The term “full” loosely refers to approaches that estimate all unknowns in the model including all parameters (irrespective of how well they are informed by the data), random effects, and predictive random variables for probabilistic interpolation at arbitrary locations (and time points in spatial-temporal data).
 Spatial modeling relies upon Gaussian processes (GPs) to model dependence and achieves superior predictive inference.
-While GPs afford flexibility and are a conspicuous choice in spatial modeling, their covariance kernels do not necessarily yield computationally exploitable structures for the covariance matrices and the computational requirements for full inference become onerous and impracticable for massive datasets.
-This presents a conundrum for statisticians: should we strive to completely retain the probabilistic rigors of theoretical inference, which may be computationally burdensome, or even impracticable, given the massive amounts of data, or should we concede uncertainty quantification and adopt algorithm-driven data analysis that would easily scale massive data sets?
+While GPs afford flexibility and are a conspicuous choice in spatial modeling, they do not generally offer computationally exploitable structures for covariance matrices.
+The computational requirements for full inference become onerous and impracticable for massive datasets.
+This presents a conundrum: should we strive to completely retain the probabilistic rigors of theoretical inference, which may be computationally burdensome, or even impracticable, given the massive amounts of data, or should we concede uncertainty quantification and adopt algorithm-driven data analysis that would easily scale massive data sets?
 Spatial data analysis at unprecedented scales requires some concessions from conventional decision-theoretic paradigms as has been reflected, for example, in a recent comprehensive case study by Zhou et al., (2022), who synthesize statistical models with machine learning algorithms to measure temporal trends and spatial distribution of housing vitality by exploiting information from multiple sources (also see the discussion by Banerjee, , 2022, for conventional Bayesian modeling perspectives and its challenges).
 Other efforts in melding machine learning methods with formal spatial statistics include deep learning for spatial data (Zammit-Mangion et al., , 2022; Wikle and Zammit-Mangion, , 2023, with the latter offering a comprehensive review of spatial machine learning) and, in a very related domain, for deep GP emulation in computer experiments (Sauer et al., 2023b, ), and spatial random forests (Georganos et al., , 2021; Talebi et al., , 2022; Saha et al., , 2023).
-Even a cursory review reveals a significant literature on statistical methods for massive spatial datasets, which is too vast to be summarized here (see, e.g., Banerjee,, 2017; Heaton et al.,, 2017, and references therein).
-Within the hierarchical setup $[data\mid process]\times[process\mid parameters]\times[parameters]$, inference proceeds from spatial processes that scale massive data sets.
-Examples range from reduced-rank processes or subsets of regression approaches (Quiñonero-Candela and Rasmussen,, 2005; Cressie and Johannesson,, 2008; Banerjee et al.,, 2008; Wikle,, 2010, and references therein), multi-resolution approaches (Nychka et al.,, 2015; Katzfuss,, 2017), and graph-based models Vecchia, (1988); Datta et al., (2016); Katzfuss and Guinness, (2021); Dey et al., (2022); Sauer et al., 2023a .
-Full inference typically requires efficient Markov chain Monte Carlo (mcmc) (Finley et al.,, 2019), variational approximations (Ren et al.,, 2011; Wu et al.,, 2022; Cao et al.,, 2023) and a significant body of literature on Gaussian Markov random field approximations (Rue et al.,, 2009; Lindgren et al.,, 2011, and references therein) in conjunction with integrated nested Laplace approximations (inla) for computing the marginal distributions of the process at given locations.
+Even a cursory review reveals a significant literature on methods for massive spatial datasets, which is too vast to be summarized here (see, e.g., Banerjee, , 2017; Heaton et al., , 2017).
+In hierarchical models $data\mid process]\times[process\mid parameters]\times[parameters]$, inference proceeds from spatial processes that scale massive data sets.
+Examples range from reduced-rank or subsets of regression approaches (see, e.g., Quiñonero-Candela and Rasmussen, , 2005; Cressie and Johannesson, , 2008; Banerjee et al., , 2008; Wikle, , 2010), multi-resolution approaches (Nychka et al., , 2015; Katzfuss, , 2017), and graph-based models Vecchia, (1988); Datta et al., (2016); Katzfuss and Guinness, (2021); Dey et al., (2022); Sauer et al., 2023a .
+Full inference typically requires Markov chain Monte Carlo (mcmc) (Finley et al., , 2019), variational approximations (Ren et al., , 2011; Wu et al., , 2022; Cao et al., , 2023) and a significant body of literature on Gaussian Markov random field approximations (Rue et al., , 2009; Lindgren et al., , 2011) in conjunction with integrated nested Laplace approximations (inla) for computing the marginal distributions of the process at given locations.
 The aforementioned methods focus upon richness of statistical inference, but almost invariably involve a significant amount of human intervention to analyze spatial data.
-Even the simplest of geostatistical datasets require exploratory data analysis to learn about aspects of the underlying process that are weakly identified by the data (Zhang,, 2004; Tang et al.,, 2021).
+Even the simplest geostatistical data require exploratory data analysis to learn about aspects of the underlying process that are weakly identified by the data (Zhang, , 2004; Tang et al., , 2021).
 Building a GeoAI system, on the other hand, will require minimizing human intervention in offering a robust framework for spatial data analysis.
@@
 The premise of this approach relies upon two basic tenets: (i) model-based statistical inference for underlying spatial processes (including multivariate processes) in a robust and largely automated manner with minimal human input; and (ii) achieving such inference for truly massive amounts of data without resorting to iterative algorithms that may require significant human intervention to diagnose convergence (such as in mcmc).
-We wish to retain the benefits of Bayesian hierarchical models, convincingly documented in the aforementioned manuscripts, with regard to spatial interpolation of the underlying stochastic process.
-In particular, we seek predictive inference over the uncountable collection of points in the spatial domain while propagating uncertainty about the model’s estimation in such inference.
+We retain the benefits of posterior learning of the underlying stochastic process.
+In particular, we obtain predictive inference over the uncountable collection of points in the spatial domain while propagating uncertainty about the model’s estimation in such inference.
 However, some compromises with regard to richness of models are inevitable from the standpoint of GeoAI.
@@
 This is possible only by “fixing” certain spatial correlation kernel parameters that are weakly identified and hinder automated inference.
-Our approach relies upon carrying out fast and exact inference based upon the matrix-variate distributions and then assimilating the inference using Bayesian predictive stacking or bps (Wolpert,, 1992; Yao et al.,, 2018).
+We execute fast and exact inference based upon the matrix-variate distributions and then assimilate the inference using Bayesian predictive stacking or bps (Wolpert, , 1992; Yao et al., , 2018).
 Stacking (Breiman, , 1996) is primarily related to predictions by minimizing the generalization error rate of one or more models.
-To analyze massive amounts of data to be accessed in a GeoAI system, we devise Bayesian transfer learning over a streaming set of spatial datasets.
-We envisage a setting where the data stream into the GeoAI system in the form of $K$ subsets, where each subset is composed of $\approx n/K$ locations with $K$ being a function of computational resources available to GeoAI.
-We analyze and transfer learning from one dataset to the next and assimilate inference from these subsets to the full data set over $n$ locations.
-The concept of delegating Bayesian calculations to a group of independent datasets is intuitive and has been explored from diverse perspectives.
-Examples include Consensus Monte Carlo (cmc) (Scott et al.,, 2016) and its kernel-based adaptations (Rendell et al.,, 2021), Bayesian meta-analysis in clinical applications (Parmigiani,, 2002, see, e.g., Chapter 4 in) and in spatially-temporally structured data (Bell et al.,, 2005; Kang et al.,, 2011); and, more recently, approaches such as “meta-kriging” using geometric medians of posterior distributions (Srivastava et al.,, 2015; Minsker et al.,, 2017; Guhaniyogi and Banerjee,, 2018, 2019; Guhaniyogi et al.,, 2019).
-However, each of these approaches have, hitherto, relied upon MCMC or other iterative algorithms for estimating the weakly identifiable parameter.
-We, instead, propose to execute transfer learning using a double stacking approach.
+Bayesian predictive stacking (bps) is a particular type of Bayesian predictive synthesis (see, e.g., Yao et al., , 2018; McAlinn and West, , 2019; McAlinn et al., , 2020; Tallman and West, , 2023; Cabel et al., , 2025, for different perspectives).
+A GeoAI system will access massive amounts of data.
+We devise Bayesian transfer learning over a streaming set of spatial datasets.
+The data streams into the GeoAI system in the form of $K$ subsets, each with $\approx n/K$ locations.
+We analyze and propagate inference from one subset to the next and assimilate inference from these subsets.
+Delegating Bayesian calculations to a group of independent datasets is intuitive and has been explored from diverse perspectives.
+Examples include Consensus Monte Carlo (cmc) (Scott et al., , 2016) and its kernel-based adaptations (Rendell et al., , 2021), Bayesian meta-analysis in clinical applications (see, e.g., Chapter 4 in Parmigiani, , 2002) and in spatially-temporally structured data (Bell et al., , 2005; Kang et al., , 2011); and, more recently, “meta” approaches using diverse distributed computing architectures (Srivastava et al., , 2015; Deisenroth and Ng, , 2015; Minsker et al., , 2017; Srivastava et al., , 2018; Guhaniyogi and Banerjee, , 2018, 2019; Guhaniyogi et al., , 2019, 2022, 2023).
+Most approaches have relied on iterative algorithms for estimating weakly identifiable parameters.
+We propose transfer learning using a double stacking approach.
 We first obtain stacked inference for each of the parameters within each subset and, then, assimilate inference across subsets by a second stacking algorithm.
@@
 Figure 1: Double Bayesian predictive stacking approach representation
-The contents of the paper evolve in accordance with Figure 1.
-Section 2 outlines Bayesian transfer learning using double bps.
-Section 3 provides details on computational aspects.
-Section 4 presents simulation experiments designed for empirically evaluating double bps.
-Section 5 analyzes a very large climate data set with observed locations in the millions at global scales.
+This article evolves according to Figure 1.
+Section 2 outlines Bayesian transfer learning using double bps (dbps), Section 3 provides details on computation, Section 4 presents simulation experiments to illustrate and evaluate double bps and its applicability to amortized inference.
+Section 5 analyzes a vegetation index data set with observed locations in the millions on global scales.
 Finally, Section 6 concludes with some discussion.
@@
 2 Bayesian transfer learning
-Transfer learning (TL) broadly refers to propagating knowledge from one task to accomplish a different task.
-We consider transferring inference from one subset to the next in a stream of subsets to analyze the entire dataset.
-Learning from spatial random fields is challenging due to inherent dependencies in the data and we seek to infer about a latent spatial process (uncountable collection of unobserved random variables) with limited data (partially observed finite realization of the process).
+Transfer learning (TL) broadly refers to propagating knowledge from one task to accomplish a different task (see Suder et al., , 2023, and references therein).
+We devise transferring inference from one subset to the next in a stream of subsets to analyze the entire dataset.
+Learning from spatial random fields is challenging because of limited information available for the latent spatial process (uncountable collection of unobserved random variables) using a partially observed finite realization.
 We develop an automated approach for GeoAI.
@@
 Spatial random fields immediately present a challenge.
-The above method delivers full inference without any loss of information only if $Y_{k}$’s are reasonably assumed to be independent (or exchangeable) across blocks.
-If each row of $Y$ corresponds to one of $n$ spatial locations so that $V$ is an $n\times n$ spatial correlation matrix, the above scheme requires that each $V_{k}$ is the spatial correlation matrix constructed from the $n_{k}$ spatial locations in $\mathscr{D}_{k}$.
-Independence across the $K$ blocks may yield reasonable inference, especially if it is possible to design the blocks such that the spatial correlation across blocks do not impact overall inference, but this will require significant human intervention or a separate AI system for designing such blocks.
+The above method delivers inference without loss of information only if $Y_{k}$’s are independent between blocks.
+If each row of $Y$ corresponds to a spatial location so $V$ is an $n\times n$ spatial correlation matrix, then each $V_{k}$ is the spatial correlation matrix constructed from $n_{k}$ spatial locations in $\mathscr{D}_{k}$.
+The independence among $K$ blocks may yield reasonable inference if we can design the blocks such that the spatial correlation between blocks does not impact the general inference.
+However, designing such blocks requires significant human intervention.
 We seek to avoid this in GeoAI.
-Instead, we devise a method for assimilating the learning from each of these blocks using predictive stacking.
+Instead, we assimilate statistical learning from each of these blocks using predictive stacking.
 We exploit the fact that $V$ is indexed by a small number of parameters in a spatial correlation kernel.
-Fixing these parameters fixes $V$, and hence $V_{k}$ for each $k$, yielding closed-form posterior inference on $\beta$ and $\Sigma$ based upon the entire dataset as described above.
+Fixing these parameters fixes $V$, and hence $V_{k}$ for each $k$, producing closed-form posterior inference on $\beta$ and $\Sigma$ based on the entire dataset as described above.
 Stacking combines these analytically accessible distributions using an optimal set of weights that are computed using a convex optimization algorithm.
 These weights are then used to reconstruct the posterior and predictive distributions for the spatial random field without imposing block independence.
-It is worth remarking here that existing “meta-kriging” and related “divide and conquer” approaches (e.g., Guhaniyogi and Banerjee,, 2018; Scott et al.,, 2016, and other references on “divide and conquer” methods provided in Section 1) analyze subsets of data using mcmc, which is expensive and not fully automated.
+Existing “meta-kriging” and related “divide and conquer” approaches (e.g., Guhaniyogi and Banerjee, , 2018; Scott et al., , 2016, and other references on “divide and conquer” methods provided in Section 1) analyze subsets of data using mcmc, which is expensive and not fully automated.
 A key distinction of the current manuscript is that we abandon all iterative estimation algorithms, let alone expensive mcmc, and focus on assimilating inference using closed forms of distributions.
-The matrix-variate formulation we pursue enhances flexibility and scope of applications.
 What remains to be resolved is the issue of fixing the parameters in $V$.
@@
 $(w_{1},\ldots,w_{J})^{{\mathrm{\scriptscriptstyle T}}}=\arg\max_{w\in S_{1}^{J}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{j=1}^{J}w_{j}p\left(Y_{i}\mid\mathscr{D}_{-i},\mathscr{M}_{j}\right)\;,$ (2.3)
-where $\mathscr{D}_{-i}$ is the dataset excluding the $i$-th observation, or block (indexed by a row) of observations in the matrix $Y$, and $\mathscr{M}=\left(\mathscr{M}_{1},\ldots,\mathscr{M}_{J}\right)$ are $J$ different models.
+where $\mathscr{D}_{-i}$ is the dataset excluding the $i$-th block (indexed by a row) of observations in $Y$, and $\mathscr{M}=\left(\mathscr{M}_{1},\ldots,\mathscr{M}_{J}\right)$ are $J$ different models.
 For any given dataset $\mathscr{D}$, these $J$ different models correspond to fixed spatial correlation kernel parameters in $V$ and deliver closed-form analytic distributions.
-Solving (2.3) minimizes the Kullback-Leibler divergence from the true predictive distribution and is executed using convex optimization (Grant and Boyd,, 2008; Fu et al.,, 2020).
-While the true predictive distribution is unknown, we can use a leave-one-out (loo) estimate of the expected value of the score (see.
-e.g., Yao et al.,, 2018, for details), which requires fitting the model $n$ times as we exclude one row of the data $\mathscr{D}$ at a time so we prefer K-fold cross-validation as a more cost-effective method for generating predictions (Breiman,, 1996).
+Solving (2.3) minimizes the Kullback-Leibler divergence from the true predictive distribution using convex optimization (Grant and Boyd, , 2008; Fu et al., , 2020).
+While the true predictive distribution is unknown, we use a leave-one-out (loo) estimate of the expected value of the score (see.
+e.g., Yao et al., , 2018, for details), which requires fitting the model $n$ times as we exclude one row of the data $\mathscr{D}$ at a time.
+We prefer K-fold cross-validation as a more cost-effective method for generating predictions (Breiman, , 1996).
 2.3 Accelerated learning for spatial random fields
-We turn to adapting the aforementioned concepts into an automated learning framework for GeoAI.
 Let $\mathcal{S}=\left\{s_{1},\ldots,s_{n}\right\}\subset\mathcal{D}$ be a set of $n$ locations yielding observations on $q$ possibly correlated outcomes collected into a $q\times 1$ vector $y(s)=(y_{1}(s),\ldots,y_{q}(s))^{{\mathrm{\scriptscriptstyle T}}}$ for each $s\in\mathcal{S}$.
@@
 Let $X=[x(s_{i})^{{\mathrm{\scriptscriptstyle T}}}]$ be $n\times p$ with rows $x(s_{i})^{{\mathrm{\scriptscriptstyle T}}}$ consisting of $p<n$ explanatory variables at location $s_{i}\in\mathcal{S}$; we assume $X$ has rank $p$.
+We introduce latent spatial processes, $\omega_{j}(s)$, for each outcome $y_{j}(s)$ to capture spatial dependence and a $q\times q$ covariance matrix, $\Sigma$, to capture non-spatial dependence among the elements of $y(s)$ within $s$.
+This matrix is typically adjusted by a scale factor $\left(\alpha^{-1}-1\right)$ to accommodate additional variation at local scales.
+For example, setting $\alpha=\sigma^{2}/\left(\sigma^{2}+\tau^{2}\right)$ where $\sigma^{2}$ and $\tau^{2}$ denote variances for the spatial process and measurement error (“nugget”) implies $\alpha$ is the ratio of the spatial variance (partial sill) to the total variance (sill) gleaned from a variogram in classical geostatistics.
 We cast this into (2.2), but explicitly introduce a latent $q\times 1$ spatial process $\omega(s)$ as
@@
 To account for measurement errors in observations, as is customary in geostatistics, we introduce a discontinuity in the spatial correlation function and modify the elements to $\rho(s_{i},s_{j};\phi)+(\alpha^{-1}-1)\mathds{1}_{s_{i}=s_{j}}$, where $\alpha\in[0,1]$ represents the proportion of total variability attributed to the spatial process.
-The covariance among the elements of $y(s)$ within $s$ is given by the elements of $\left(\alpha^{-1}-1\right)\Sigma$.
-Thus, $\Sigma$ is the within-location (nonspatial) dependence among the outcomes adjusted by a scale of $\left(\alpha^{-1}-1\right)$ to accommodate additional variation at local scales.
-The parameter $\alpha$ is the ratio of the “partial sill” to the “sill” in classical geostatistics.
-In the special case when $\Sigma=\sigma^{2}\mathbb{I}_{q}$, we interpret $\alpha=\sigma^{2}/\left(\sigma^{2}+\tau^{2}\right)$, where $\tau^{2}$ is the measurement error variance (“nugget”) and $\alpha$ is the ratio of the spatial variance (partial sill) to the total variance (sill).
 Letting $\gamma^{{\mathrm{\scriptscriptstyle T}}}=\left[\beta^{{\mathrm{\scriptscriptstyle T}}},\Omega^{{\mathrm{\scriptscriptstyle T}}}\right]$ be $q\times(p+n)$, we assume $\{\gamma,\Sigma\}\sim\operatorname{MNIW}\left(\mu_{\gamma},V_{\gamma},\Psi_{0},\nu_{0}\right)$, where
@@
 $p(\gamma,\Sigma\mid\mathscr{D})=\operatorname{MNIW}\left(\gamma,\Sigma\mid\mu_{\gamma}^{\star},V^{\star}_{\gamma},\Psi^{\star},\nu^{\star}\right),$ (2.6)
-where $V^{\star}_{\gamma}=\left[\begin{array}[]{cc}\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}X+M_{0}^{-1}&\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}\\ \frac{\alpha}{1-\alpha}X&\rho_{\phi}^{-1}(\mathcal{S},\mathcal{S})+\frac{\alpha}{1-\alpha}\mathbb{I}_{n}\end{array}\right]^{-1}$ and $\mu_{\gamma}^{\star}=V^{\star}_{\gamma}\left[\begin{array}[]{c}\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}Y+m_{0}\\ \frac{\alpha}{1-\alpha}Y\end{array}\right]$, $\Psi^{\star}=\Psi_{0}+\frac{\alpha}{1-\alpha}Y^{{\mathrm{\scriptscriptstyle T}}}Y+m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0}m_{0}-\mu_{\gamma}^{*{\mathrm{\scriptscriptstyle T}}}V^{\star-1}_{\gamma}\mu_{\gamma}^{\star}$ and $\nu^{\star}=\nu_{0}+n$.
+where $V^{\star}_{\gamma}=\begin{bmatrix}\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}X+M_{0}^{-1}&\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}\\ \frac{\alpha}{1-\alpha}X&\rho_{\phi}^{-1}(\mathcal{S},\mathcal{S})+\frac{\alpha}{1-\alpha}\mathbb{I}_{n}\end{bmatrix}^{-1}$ and $\mu_{\gamma}^{\star}=V^{\star}_{\gamma}\begin{bmatrix}\frac{\alpha}{1-\alpha}X^{{\mathrm{\scriptscriptstyle T}}}Y+m_{0}\\ \frac{\alpha}{1-\alpha}Y\end{bmatrix}$, $\Psi^{\star}=\Psi_{0}+\frac{\alpha}{1-\alpha}Y^{{\mathrm{\scriptscriptstyle T}}}Y+m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0}m_{0}-\mu_{\gamma}^{*{\mathrm{\scriptscriptstyle T}}}V^{\star-1}_{\gamma}\mu_{\gamma}^{\star}$ and $\nu^{\star}=\nu_{0}+n$.
 The framework in (2.4) is equivalent to (2.2) with $V=R_{\phi}+(\alpha^{-1}-1)\mathbb{I}_{n}$ with $R_{\phi}=[\rho(s_{i},s_{j};\phi)]$.
@@
 where $\nu^{\star}_{[-l]}=\nu_{0}+n_{k,[-l]}$, $n_{k,[-l]}$ is the cardinality of $\mathcal{S}_{k,[-l]}$ (which is the set of locations in $\mathscr{D}_{k,[-l]}$), and $\Psi^{\star}_{[-l]}=\Psi_{0}+(\alpha_{j}^{-1}-1)Y^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}^{{\mathrm{\scriptscriptstyle T}}}M_{0}m_{0}-\mu^{{\mathrm{\scriptscriptstyle T}}\star}_{\gamma,[-l]}V^{-1\star}_{[-l]}\mu^{\star}_{\gamma,[-l]}$.
-While $V^{\star}_{i}=M_{y,i}V_{\gamma}^{\star}M_{y,i}^{{\mathrm{\scriptscriptstyle T}}}+V_{\Omega_{i}}+(\alpha_{j}^{-1}-1)$, $\mu^{\star}_{i}=M_{y,i}\mu^{\star}_{\gamma,[-l]}$, which are defined by computing the following auxiliary quantities: $V^{-1\star}_{\gamma,[-l]}=\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}X_{k,[-l]}+M_{0}^{-1}&\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}\\ \frac{\alpha_{j}}{1-\alpha_{j}}X_{k,[-l]}&\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})+\frac{\alpha_{j}}{1-\alpha_{j}}\mathbb{I}_{n_{k,[-l]}}\end{bmatrix}$, $\mu^{\star}_{\gamma,[-l]}=V^{\star}_{\gamma,[-l]}\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}\\ \frac{\alpha_{j}}{1-\alpha_{j}}Y_{k,[-l]}\end{bmatrix}$, $M_{y,i}=\begin{bmatrix}X_{k,i}&\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})\end{bmatrix}$, and $V_{\Omega_{i}}=\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,i})-\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})\rho_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,i})$.
+While $V^{\star}_{i}=M_{y,i}V_{\gamma}^{\star}M_{y,i}^{{\mathrm{\scriptscriptstyle T}}}+V_{\Omega_{i}}+(\alpha_{j}^{-1}-1)$, $\mu^{\star}_{i}=M_{y,i}\mu^{\star}_{\gamma,[-l]}$, which are defined by computing the following auxiliary quantities: $V^{-1\star}_{\gamma,[-l]}=\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}X_{k,[-l]}+M_{0}^{-1}&\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}\\ \frac{\alpha_{j}}{1-\alpha_{j}}X_{k,[-l]}&\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})+\frac{\alpha_{j}}{1-\alpha_{j}}\mathbb{I}_{n_{k,[-l]}}\end{bmatrix}$, $\mu^{\star}_{\gamma,[-l]}=V^{\star}_{\gamma,[-l]}\begin{bmatrix}\frac{\alpha_{j}}{1-\alpha_{j}}X^{{\mathrm{\scriptscriptstyle T}}}_{k,[-l]}Y_{k,[-l]}+m_{0}\\ \frac{\alpha_{j}}{1-\alpha_{j}}Y_{k,[-l]}\end{bmatrix}$, $M_{y,i}=\begin{bmatrix}X_{k,i}&\quad\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})\end{bmatrix}$, and $V_{\Omega_{i}}=\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,i})-\rho_{\phi_{j}}(\mathcal{S}_{k,i},\mathcal{S}_{k,[-l]})\rho^{-1}_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,[-l]})\rho_{\phi_{j}}(\mathcal{S}_{k,[-l]},\mathcal{S}_{k,i})$.
 Note that $\nu^{\star}_{[-l]}$ is scalar, $\mu^{\star}_{i}$ is $(1\times q)$ row vector, $V^{\star}_{i}$ is a scalar, and $\Psi^{\star}_{[-l]}$ is $(q\times q)$ matrix.
 Further details, including derivations and implementation, are supplied in Section LABEL:sec:matrixT and Algorithm 1.
-For each dataset $\mathscr{D}_{k}$, bps computes: (i) an estimate of the posterior predictive $\hat{p}(\cdot\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$ for $k=1,\dots,K$ and $j=1,\dots,J$; and (ii) a set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$.
-Once these weights are available, we apply bps a second time to obtain a weighted average of $\hat{p}(\cdot\mid\mathscr{D}_{k})$ over the $k$ subsets.
-This double bps of predictive densities seeks weights $w=\{w_{k}\}_{k=1,\ldots,K}$ such that $\hat{w}=\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}\hat{p}\left(Y_{i}\mid\mathscr{D}_{k}\right)$ (see Section LABEL:sec:BPSobjfun).
-Once the stacking weights $\hat{w}=\{\hat{w}_{k}\}_{k=1,\dots,K}$ are obtained, the double bps estimation of any posterior or posterior predictive distributions of interest is obtained from
-$\hat{p}(\cdot\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)\;.$ (2.10)
+For each dataset $\mathscr{D}_{k}$, bps computes: (i) an estimate of the posterior predictive $\hat{p}(\,\cdot\;;\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\,\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$ for $k=1,\dots,K$ and $j=1,\dots,J$; and (ii) a set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$.Once these weights are available, we apply bps a second time to obtain a weighted average of $\hat{p}(\,\cdot\;;\mathscr{D}_{k})$ over the $k$ subsets.
+This double bps (dbps) of predictive densities seeks weights $w=\{w_{k}\}_{k=1,\ldots,K}$ such that $\hat{w}=\max_{w\in S_{1}^{K}}\frac{1}{n}\sum_{i=1}^{n}\log\sum_{k=1}^{K}w_{k}\hat{p}\left(Y_{i}\mid\mathscr{D}_{k}\right)$ (see Section LABEL:sec:BPSobjfun).
+Once stacking weights $\hat{w}=\{\hat{w}_{k}\}_{k=1,\dots,K}$ are computed using Algorithm 2, sampling from the posterior and posterior predictive distributions is obtained from
+$\hat{p}(\,\cdot\;;\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\,\cdot\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)\;.$ (2.10)
 Only posterior predictive distributions are considered to acquire the two sets of stacking weights.
@@
 This makes sampling from (2.10) straightforward.
-First, the set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$ obtained using bps within the subset of the data $\mathscr{D}_{k}$, is primarily used to approximate the subset posterior distribution $\hat{p}(\Theta\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}p\left(\Theta\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$, for each subset $k=1,\dots,K$.
-By considering the second set of stacking weights $\hat{w}=\{\hat{w_{k}}\}_{k=1,\dots,K}$, the stacked full posterior distribution is $\hat{p}(\Theta\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\hat{p}(\Theta\mid\mathscr{D}_{k})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p(\Theta\mid\mathscr{D}_{k},\mathscr{M}_{j})$, where $\Theta=\{\gamma,\Sigma\}$.
+First, the set of stacking weights $\hat{z}_{k}=\{\hat{z}_{k,j}\}_{j=1,\dots,J}$ obtained using bps within the subset of the data $\mathscr{D}_{k}$, is primarily used to approximate the subset posterior distribution $\hat{p}(\Theta\;;\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}p\left(\Theta\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$, for each subset $k=1,\dots,K$.
+By considering the second set of stacking weights $\hat{w}=\{\hat{w_{k}}\}_{k=1,\dots,K}$, the stacked full posterior distribution is $\hat{p}(\Theta\;;\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\hat{p}(\Theta\;;\mathscr{D}_{k})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p(\Theta\mid\mathscr{D}_{k},\mathscr{M}_{j})$, where $\Theta=\{\gamma,\Sigma\}$.
 This is a substantial simplification over meta-kriging (Guhaniyogi and Banerjee, , 2018, 2019) that does not require any empirical approximation of posterior or predictive distributions.
 Again, the predictive random variable $Y_{\mathcal{U}}$ is recovered from (2.10) as
-$\hat{p}(Y_{\mathcal{U}}\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(Y_{\mathcal{U}}\mid\mathscr{D}_{k},\mathscr{M}_{j}\right).$ (2.11)
+$\hat{p}(Y_{\mathcal{U}}\;;\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(Y_{\mathcal{U}}\mid\mathscr{D}_{k},\mathscr{M}_{j}\right).$ (2.11)
 Inferential interest resides with the posterior predictive surface for $\Omega$.
-Thus, we estimate $\Omega_{\mathcal{U}}$ using $\hat{p}(\Omega_{\mathcal{U}}\mid\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\hat{p}(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k})$, where $\hat{p}(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$.
+Thus, we estimate $\Omega_{\mathcal{U}}$ using $\hat{p}(\Omega_{\mathcal{U}}\;;\mathscr{D})=\sum_{k=1}^{K}\hat{w}_{k}\;\hat{p}(\Omega_{\mathcal{U}}\;;\mathscr{D}_{k})$, where $\hat{p}(\Omega_{\mathcal{U}}\;;\mathscr{D}_{k})=\sum_{j=1}^{J}\hat{z}_{k,j}\;p\left(\Omega_{\mathcal{U}}\mid\mathscr{D}_{k},\mathscr{M}_{j}\right)$.
 Algorithm 1 Computing stacking weights within subsets using bps
@@
 25:return $\{\hat{z},\{pd_{k,j,i}\},G_{all}\}$
-3 Computer programs and resources
-Our subsequent implementations, including all simulations experiments and data analysis, are executed in native R and c++ deploying the spBPS package.
-In particular, all programs used to obtain the ensuing analysis are publicly accessible from the GitHub repository lucapresicce/Bayesian-Transfer-Learning-for-GeoAI that links the Rcpp-based spBPS package including the main function executing bps for univariate and multivariate geostatistical models.
-The reported results are executed on a standard laptop, which runs an Intel Core I7-$8750$H CPU with $5$ available cores for parallel computation and $16$ GB of RAM.
-We conduct experiments comparing double bps with spatial meta-kriging (smk) using a meta-posterior (Minsker et al.,, 2017) defined as the geometric median of individual posterior distributions based on subsets of the data (as implemented in Guhaniyogi and Banerjee,, 2018, 2019).
-We execute meta-kriging using the spMvLM function within the spBayes R package that fits a linear model of coregionalization for multivariate spatial data using MCMC (Finley et al.,, 2015).
-For parallel implementations of bps and smk, we employ R packages doParallel, and foreach (Corporation and Weston,, 2022; Microsoft and Weston,, 2022).
-We map the interpolated spatial surfaces using the MBA R package (Finley et al.,, 2022), while sampling from the matrix-variate normal and t distributions are achieved using the mvnfast R package (Fasiolo,, 2014).
-Section LABEL:sec:compdetails in the Supplement provides further details including memory management considerations and efficient solutions to spatial “BIG” data analysis, and sensitivity on the number of data shards $K$.
 Algorithm 2 Calculating stacking weights between subsets using bps
@@
 6:return $\hat{w}=\{\hat{w}_{k}:k\in\{1,\dots,K\}\}$
+3 Computer programs and resources
+All our subsequent analysis are implemented in native R and c++ deploying the spBPS package.
+All programs required to reproduce the analysis are publicly accessible from the GitHub repository lucapresicce/Bayesian-Transfer-Learning-for-GeoAI that links the Rcpp-based spBPS package.
+The reported results are from a standard laptop running an Intel Core I7-87508750H CPU with $5$ cores for parallel computation and $16$ GB of RAM.
+We fit a linear model of coregionalization (lmc, Banerjee et al., , 2015) and multivariate seemingly unrelated Bayesian additive regression trees (multivariate bart Esser et al., , 2025) using spBayes and suBART packages, respectively.
+We also compare with machine learning methods and AI systems using a scalable platform for parallelized supervised and unsupervised machine learning algorithms offered by h2o (Fryda et al., , 2024).
+We specifically fit distributed random forest (drf), gradient boosting (gbm), deep neural network (dnn), and a fully automatic machine learning algorithm (automl).
+For parallel implementations of dbps, we employ R packages doParallel, and foreach (Corporation and Weston, , 2022; Microsoft and Weston, , 2022).
+We map the interpolated spatial surfaces using MBA (Finley et al., , 2022), while sampling from the matrix-variate normal and t distributions is achieved using mvnfast (Fasiolo, , 2014).
+Section LABEL:sec:compdetails specifies computational considerations and sensitivity to the number of data shards, $K$, for spatial “BIG” data analysis.
+We build a Bayesian transfer learning engine to conduct amortized Bayesian inference (Zammit-Mangion et al., , 2024) using dbps.
+We implement a residual neural network (ResNet) (He et al., , 2015) using the R interfaces supplied by tensorflow (Allaire et al., , 2024) and keras (Kalinowski et al., , 2024) for native Python.
 4 Simulation experiments
-We conduct multiple simulation experiments to evaluate inferential performance of double bps while underscoring comparisons with smk.
-We present a selection of simulation results for multivariate models here and refer the reader to Section LABEL:sec:sim_suppl for further experiments.
-4.1 Simulation - BPS vs SMK
-In the first setting, we consider two synthetic datasets with common structures but distinct sizes using the model (2.4).
-Both datasets consist of $p=2$ predictors and $q=2$ response variables, but different numbers of spatial locations, $n=5,000$ and $n=10,000$, respectively.
-We generate spatial coordinates from a uniform distribution on $[0,1]^{2}$ domain and fix them.
-We build the $n\times n$ spatial correlation matrix $V$ over these coordinates using an exponential spatial correlation function with $\phi=4$ and specify $\Sigma=\mathbb{I}_{q}$.
-From these specifications we generate the $n\times q$ matrix $Y$ from the first equation of (2.4) with fixed $p\times q$ matrix $\beta=\begin{bmatrix}-0.75&1.85\\ 0.90&-1.10\end{bmatrix}$, a fixed $n\times p$ matrix $X$ with a first column of ones, representing the intercept, and $p-1$ columns of values randomly simulated from a uniform distribution on $[0,1]$ (emulating standardized predictors), and the proportion of spatial variability $\alpha=0.8$.
-We analyze the simulated datasets using double bps and smk in two settings: (i) $K=10$ and $n=5,000$; (ii) $K=10$ and $n=10,000$; (iii) $K=5$ and $n=5,000$; and (iv) $K=20$ and $n=10,000$.
-These settings produce subsets of size $n/K\in\{500,1000\}$.
-We implement bps using $J=9$ candidate models $\mathscr{M}_{j},j=1,\dots,J$, where each model is specified by a set of candidate values for the hyperparameters ${\alpha_{j},\phi_{j}}$ in (2.4).
-These hyperparameters represent the proportion of spatial variability and the parameter(s) of the spatial correlation function, respectively.
-The set of candidate models is constructed as the set of all possible combinations of values for these hyperparameters.
-In the subsequent experiments, the grid of models was built using $\alpha\in\{0.70,0.80,0.90\}$ and $\phi\in\{3,4,5\}$.
-These values resemble an effective spatial range of $\{0.99,0.75,0.60\}$ units, corresponding to $70\%,53\%,42\%$ of the maximum inter-site distance inside the unit square, beyond which the spatial correlation drops below $0.05$.
-Equation (2.5) follows choices in Zhang et al., (2021).
-Specifically, we set $m_{0}=0_{p\times q}$, $M_{0}=10\mathbb{I}_{p}$, $\Psi_{0}=\mathbb{I}_{q}$, and $\nu_{0}=3$ in the $\operatorname{MNIW}$ joint prior for $\{\gamma,\Sigma\}$ in 2.5.
-For data analysis, we use an exponential spatial correlation function for $\rho_{\phi}(\cdot,\cdot)$, which is completely defined conditionally on $\mathscr{M}_{j}$, since specifies a value for $\phi$.
-Again, in the conjugate framework, we draw $R=250$ posterior samples used for inference.
-Model $n=5000,K=10$ $n=5000,K=5$ $n=10000,K=20$ $n=10000,K=10$
-smk 51.41 237.23 103.36 446.01
-bps 1.38 7.22 2.24 10.58
-Relative ratio 37.25 32.85 46.14 42.15
-Table 1: Computational aspects for the different settings, with time in minutes.
-We also apply smk to the two simulated datasets with the same combinations of $n$ and $K$ as bps.
-Unlike bps, where we stack analytically tractable posteriors over a range of fixed values of spatial covariance kernel parameters, the smk implementation attempts full Bayesian inference using prior distributions on spatial covariance kernel parameters.
-We fit the linear model of coregionalization described in Finley et al., (2015) for each subset of the multivariate spatial data using MCMC.
-The posterior samples from the $K$ subsets are combined using Weiszfeld’s iterative algorithm (Minsker et al.,, 2017) to produce an estimate of the geometric median of the posterior distributions.
-Table 1, which compares the computational speed of double bps with smk reveals massive computational gains accrued from bps.
-The computational advantage evinced from the relative ratio become more pronounced as the size of the data becomes larger despite the larger subsets.
-This is explained by the fact that fitting the Gaussian process regression dominates the computation relative to the assimilation of inference from the subsets.
-If the number of locations explode, then the geometric median of posteriors required by smk is computationally unfeasible.
-While smk offers Bayesian estimates using MCMC for each subset, double bps avoids MCMC and, hence, issues of convergence.
-Importantly, as we see in Figure 2, which depicts estimated response surfaces using bps and smk corresponding to $n=5000$ and $K=10$, inference from smk and bps are practically indistinguishable.
-The root mean squared prediction errors, rmspe, reported in Figure 2 denotes the average squared differences between the generated and estimated values of the response and reveals minor discrepancies between bps and smk.
-Section LABEL:sec:sim_suppl_M presents results for the other configurations of $n$ and $K$, each of which reveals that bps offers practically indistinguishable spatial interpolation from smk at a fraction of the computational cost.
-Figure 2: from left to right: comparison between the true generated response surfaces, the surfaces predicted from bps and smk (posterior mean), with rmspe. For $n=5000$, $K=10$.
-Figure 3 reports 95% posterior predictive intervals for the response.
-Again, the empirical coverage is impressive.
-While we see slightly wider intervals from BPS, this is less pronounced than the underestimation of variability seen with smk.
-Moreover, Figure 3 reveals superior map estimates for the bps.
-Finally, Figure 4 presents the recovery of parameter estimates.
-As seen in predictive inference, the posterior credible intervals for parameters also deliver practically indistinguishable inference for the two modeling frameworks.
-In particular, both methods recover the true values for $\beta$ and $\Sigma$, while BPS reconstructs a better point estimate for range parameters $\phi$ using $\sum_{k=1}^{K}\hat{w}_{k}\sum_{j=1}^{J}\hat{z}_{j}\phi_{j}$.
-Similar conclusions are drawn from other values of $n$ and $K$ explored in Section LABEL:sec:sim_suppl for univariate responses.
-Figure 3: from top to bottom: comparison between posterior predictive intervals for the predicted response from bps and smk, with empirical coverage. For $n=5000$, $K=10$.
-Figure 4: from left to right: comparison between posterior credible intervals for the parameters recovered from bps and smk. For $n=5000$, $K=10$.
-4.2 Simulation - Multivariate Transfer learning
-We now consider inferential performance for different misspecified models.
-As in the earlier example, we generate the $n\times q$ outcome $Y$ using  (2.4) with $n=5,000$, $q=3$ and $p=2$, $X$ includes an intercept and $p-1$ columns generated from a standard uniform distribution over $[0,1]$, $\beta=\begin{bmatrix}-0.75&1.05&-0.35\\ 2.20&-1.10&0.45\end{bmatrix}$ and $\Sigma=\begin{bmatrix}2.00&0.80&0.20\\ 0.80&2.00&-0.45\\ 0.20&-0.45&2.00\end{bmatrix}$.
-The $n\times n$ spatial correlation matrix $V$ is specified using an exponential correlation matrix with a fixed range parameter at $\phi=4$ and $\alpha=0.8$ to be the strength of spatial dependence.
-We analyze the generated data using three different models that assign prior distributions to $\{\beta,\Sigma\}$ with $m_{0}=0_{p\times q}$, $M_{0}=10\mathbb{I}_{p}$, $\Psi_{0}=\mathbb{I}_{q}$, and $\nu_{0}=3$, but different values of $\alpha$ and $\phi$ representing (i) well-specified setting with the data generating values $\{\alpha=0.8,\phi=4\}$; (ii) moderately misspecified setting with $\{\alpha=0.45,\phi=6.63\}$; and (iii) a highly misspecified setting with $\{\alpha=0.25,\phi=50\}$.
-For each setting, we partition the original data into $K=10$ subsets, corresponding to subsets of size $500$.
-We consider $J=9$ competitive models for bps and consider $\alpha\in\{0.75,0.80,0.85\}$ and $\phi\in\{2,4,6\}$ that yields effective spatial ranges in the percentage of maximum point inter-distance of $105\%,53\%,35\%$ respectively.
-Here, we display results for the bps, while graphical results for other settings are available in Section LABEL:sec:sim_suppl_M.
-Figure 5 presents a visual comparison between the generated true response and the posterior mean of spatial predictions delivered by bps.
-We present the rmspe along with a set of credible intervals for the parameters relative to the true values.
-The inferential performance evaluated using the rmspe as well as the spatially interpolated maps corroborates the effectiveness of bps.
-Figure 6 reveals that more than half of the model unknowns are included in the 95% posterior credible interval.
+We evaluate computational and inferential performances of double bps, while underscoring comparisons with multiple alternative methodologies.
+We present a selection of simulation results for multivariate models here and refer the reader to Section LABEL:sec:sim_suppl_M for further experiments.
+4.1 Simulation - Transfer learning in $\mathscr{M}$-closed & $\mathscr{M}$-open settings
+We evaluate inferential and predictive performance under different settings.
+We explore how double bps behaves in the $\mathscr{M}$-closed and $\mathscr{M}$-open settings and compare with an exact transfer learning framework we devise in Section 2.1.
+In the latter, model specification is characterized by different values of $\alpha$ and $\phi$ representing (i) well-specified (ws) setting with the data generating values $\{\alpha=0.8,\phi=4\}$; (ii) moderately misspecified (ms) setting with $\{\alpha=0.45,\phi=6.63\}$; and (iii) a highly misspecified (hms) setting with $\{\alpha=0.25,\phi=50\}$.
+The double bps was tested over $\mathscr{M}$-closed and $\mathscr{M}$-open settings.
+The former considers situations where the true model exists, and it is identified within a finite set of considered models.
+Here, the “true” model is the one such that $\{\alpha=0.8,\phi=4\}$.
+Then, for double bps under closed setting (bps-c) we specify $J=9$ competitive models with $\alpha\in\{0.75,0.80,0.85\}$ and $\phi\in\{2,4,6\}$ that yields effective spatial ranges in the percentage of maximum point inter-distance of $105\%,53\%,35\%$ respectively, including the true model as one of the possible candidates.
+Conversely, in the $\mathscr{M}$-closed setting, even though the true model exists, it cannot be fully specified.
+Thus, for double bps under open setting (bps-o), we randomly define $J=9$ candidate models.
+In particular, we uniformly sample 3 values for $\alpha\in(0,1)$ and 3 values for $\phi\in(0,50)$.
+We perform the experiment using 50 replications.
+Each replicate consists of values of $n\times q$ outcome $Y$ generated from (2.4) with $n=5,000$, $q=3$ and $p=2$, $X$ includes an intercept and a predictor generated from a standard uniform distribution over $[0,1]$, $\beta=\begin{bmatrix}-0.75&1.05&-0.35\\ 2.20&-1.10&0.45\end{bmatrix}$ and $\Sigma=\begin{bmatrix}2.00&0.80&0.20\\ 0.80&2.00&-0.45\\ 0.20&-0.45&2.00\end{bmatrix}$.
+The $n\times n$ spatial correlation matrix $V$ is specified using an exponential correlation function with $\phi=4$ and $\alpha=0.8$.
+Figure 2 presents (i) mean square prediction error (mspe); (ii) predictive interval width; (iii) absolute bias; and (iv) variance.
+We present boxplots for the distribution of each metric over the 50 replicates.
+This is made for each response and for each setting.
+In terms of predictive mspe, absolute bias, and variance, the settings bps-c, bps-o depict sligthly better performance.
+However, this seems to be compromising predictive interval width as the uncertainty of prediction results is much higher.
+Approaches that estimate $\{\alpha,\phi\}$, rather than fix them, more uncertainty is somehow expected.
+In addition, we find no evidence of any difference between bps-c, and bps-o for any metric.
+This is surprising, as it suggests the reliability of dbps even under the transfer learning setting we devise in Section 2.3.
+Finally, irrespective of $\mathscr{M}$-closed or $\mathscr{M}$-open settings, it is more convenient to specify a set of candidate models using double bps instead of attempting to fix $\{\alpha,\phi\}$.
+Figure 3 presents posterior inference for (i) average empirical bias; (ii) average coverage; and (iii) average standard deviation, where the average is taken over the 50 replications.
+As expected, misspecification induces empirical bias in posterior estimates.
+The top panel reports how greater levels of bias are associated with models that are farther away from the truth.
+Here, double bps is placed in the middle for both $\mathscr{M}$ settings.
+The middle plot in Figure 3 shows that $\Sigma$ is the most affected parameter by misspecifications.
+Its elements are well captured only by double bps, with coverage close to nominal, followed by ws specification, which performs worse.
+A Similar pattern holds for $\beta$, where only double bps ensures adequate coverage.
+This reflects greater posterior variability in stacking approaches, as shown in the bottom panel reporting posterior standard deviations across settings.
 Unsurprisingly, inferential performance is weakest for the spatial variance, which is not identifiable from the realized data (Zhang, , 2004).
-Further results on simulation experiments are available in Section LABEL:sec:sim_suppl_M.
-Figure 5: from top to bottom: comparison between the true generated spatial surfaces and the posterior mean surfaces of predictions from bps, with rmspe for each outcome.
-Figure 6: Posterior credible intervals for parameters recovered by bps, with true parameters in yellow.
-Table 2 presents predictive performance in terms of rmspe, along with the hyperparameter values corresponding to the different experimental settings.
-The first three rows correspond to fixed values of $\{\alpha,\phi\}$ in the Bayesian transfer learning approach of Section 2.1.
-The last row corresponds to bps when $\alpha$ and $\phi$ are weighted averages using the stacking weights as described in the last paragraph of Section 4.1.
-While bps outperforms transfer learning in terms of rmspe, some remarks on transfer learning are merited.
-The transfer learning approach offers slightly better predictive performance than the well-specified model with misspecified data than with well-specified data.
-This, while somewhat unexpected, is attributed to the estimated response surfaces being practically indistinguishable.
-Furthermore, transfer learning in even the highly misspecified setting performs competitively.
-In conclusion, these findings seem to be consistent with theoretical insights offered by Stein, (1988) and Stein and Handcock, (1989), who theoretically established the ability of Gaussian processes to deliver good predictive performance even for misspecified covariance functions within the in-fill asymptotic paradigm.
-Model $\alpha$ $\phi$ Average rmspe
-well specified 0.8 4 0.942
-misspecified 0.45 6.63 0.963
-highly misspecified 0.25 50 1.236
-double bps 0.75 4.11 0.900
-Table 2: Average rmspe and specification for each model scenario.
-5 Application
-Global warming and its critical consequences (see Fisher,, 1958; Nicholls,, 1989; Friehe et al.,, 1991; O’Carroll et al.,, 2019) are increasingly being investigated using machine learning and AI tools that assist in developing predictive global models and formulating data-driven policies.
-trying to produce faster solutions.
-The key aspects of global warming are related to global temperature, atmospheric and land surface compositions.
-We analyze data extracted from NASA’s Moderate Resolution Imaging Spectroradiometer (modis) platform.
-Specifically, we consider as response variables two vegetation indices present in vi data: the normalized difference vegetation index (ndvi); and the red reflectance (rr).
-Both measure the activity of vegetation on land surfaces and are commonly investigated to help understand biophysical and structural properties of vegetation at global scales (Walther et al.,, 2018; Jiang et al.,, 2020; Ai et al.,, 2020).
+double bps behaves very similarly among $\mathscr{M}$-closed and $\mathscr{M}$-open.
+Additional simulation experiments are discussed in Section LABEL:sec:sim_suppl_M.
+Figure 2: Predictive MSPE, interval width, absolute bias, and variance boxplot across responses and settings from 50 replications.
+Figure 3: Average posterior bias, coverage, and standard deviation across parameters and settings from 50 replications.
+While dbps outperforms exact transfer learning (Section 2) in $\mathscr{M}$-closed and $\mathscr{M}$-open settings, the latter performs competitively with improved predictive performance in ws and ms settings than with hms.
+Our overall findings appear consistent with theoretical insights that Gaussian processes tend to deliver good predictive performance even for misspecified covariance functions in fixed domains (Stein, , 1988; Stein and Handcock, , 1989).
+4.2 Amortized Bayesian inference
+We conduct transfer learning by supervising a neural network using the outputs of double bps to deliver amortized Bayesian inference.
+We generate $100$ instances of $Y$ from (2.4) using a fixed realization of $\Omega$ for $q=2$ correlated outcomes, $n=500$ spatial locations that remain fixed across the datasets, and a fixed design matrix $X$ with $p=2$ comprising an intercept and a single predictor whose values were sampled independently from a uniform distribution over $[0,1]$.
+The true regression coefficients are fixed at $\beta=\begin{bmatrix}-0.75&1.85\\ 0.9&-1.10\end{bmatrix}$, with $\Sigma=\begin{bmatrix}1&-0.3\\ -0.3&1\end{bmatrix}$, $\alpha=0.8$, and $\rho_{\phi}(s_{i},s_{j})=\exp(-\phi\|s_{i}-s_{j}\|)$ with $\phi=4$.
+We train the neural network using $R=250$ posterior samples by applying double bps to each generated dataset with $K=5$ subsets, $\alpha\in\{0.7,0.8,0.9\}$ and $\phi\in\{3,4,5\}$.
+These yield $100$ instances of $\{Z,\Theta\}$, where $Z=[Y:X]\in\mathbb{R}^{n\times(q+p)}$ and $\Theta\in\mathbb{R}^{[(qp)+(q(q+1)/2)+(nq)]\times 3}$ comprises the $\{2.5,50,97.5\}$ posterior quantiles for the distinct elements of $\{\beta,\Sigma,\Omega\}$.
+We use a deep neural network comprising 3 hidden layers with 128, 256, and 512 nodes, with ReLU activations.
+The residual network is trained over 50 epochs with 24 batches per epoch.
+For evaluation, we apply the trained model to unseen datasets with the same dimensions.
+Figure 4 displays the predicted posterior intervals (blue bars) for $\{\beta,\Sigma\}$, along with the true values of the parameters (yellow triangles).
+The credible intervals $95\%$ capture all true values and the predicted medians ($50\%$) align closely with the true values, highlighting the strong predictive precision of the deep network in recovering posterior summaries.
+Figure 5 displays the predicted posteriors for $\Omega$ quantiles.
+We compare the results from amortized inference, with the true $\Omega$, and the double bps prediction for $50$th quantile, presented in the first and the second column of Figure 5, respectively.
+These results illustrate the strengths of amortized inference and transfer learning.
+Once trained, the deep network provides instant posterior quantile estimates for new datasets, without requiring us to rerun double bps.
+This amortizes the computational cost for future tasks.
+Additionally, the model generalizes across a range of data-generating conditions effectively enabling posterior transfer learning to new but structurally similar problems.
+This makes the approach especially useful in large-scale or resource-constrained applications where repeated full Bayesian inference is prohibitively expensive.
+Figure 4: Amortized posterior credible intervals for parameters. True parameters in yellow.
+Figure 5: Surface interpolations for true spatial process, BPS prediction (50 quantile), and Amortized prediction of $\{50,2.5,97.5\}$ quantiles. Each row corresponds to an outcome.
+5 Data analysis
+5.1 Vegetation index data
+Statisticians and scientists face growing demands to analyze and study global warming datasets (see, e.g., Fisher, , 1958; Nicholls, , 1989; Friehe et al., , 1991; O’Carroll et al., , 2019).
+The sheer volume of such datasets and the process-driven models required for their analysis have naturally raised the question of migrating such analysis to AI platforms.
+Our current application analyzes vegetation index data from the Moderate Resolution Imaging Spectroradiometer (modis) developed by nasa.
+We specifically focus on “mod13c1.061 - Terra Vegetation Indices 16-Day L3 Global 0.05 Deg Climate Modeling Grid” (Didan, , 2021), which provides vegetation indices per pixel on a 0.05-degree climate modeling grid (3600 rows by 7200 columns of 5600-meter pixels).
+It contains 16-day global composites, cloud-contamination-free, with additional reflectance and angular information.
+Modeling Normalized Difference Vegetation Index (ndvi) and Red Reflection (rr) jointly is scientifically important because it separates and defines the specific biophysical factors influencing vegetation.
+ndvi is built on the fundamental principle that healthy vegetation absorbs red light and reflects near-infrared (NIR) light.
+By examining both the calculated NDVI and its red reflectance component, scientists can gain deeper insights into vegetation’s structural and biochemical properties.
+The solar zenith angle (sza), which is the angle between the Sun and the point directly overhead, is a shared predictor that influences how much solar irradiation reaches the surface and is, therefore, crucial when assessing biomass and vegetation indices.
+ndvi and rr reflect the ability of vegetation to absorb photosynthetically active radiation and help scientists understand processes underlying climate change (Tucker, , 1979; Sellers, , 1985; Justice et al., , 1985; Haque et al., , 2024).
+Vegetation index Mean Std.Dev Min. Max. Histogram Boxplot
+ndvi 8.593 0.517 6.909 9.469 [Uncaptioned image] [Uncaptioned image]
+Red Reflectance 8.563 0.447 8.007 9.472 [Uncaptioned image] [Uncaptioned image]
+Table 1: Summary statistics and visual representation of response variables.
+The dataset comprises $1,500,000$ locations.
+We randomly selected $1,002,500$ sites, reserving $2,500$ for predictive evaluation.
+Following Zhang and Banerjee, (2022), outcomes were log-transformed and we labeled $\log(\textsc{ndvi}+1)$ as ndvi and $\log(\textsc{rr}+1)$ as rr.
+All variables, including sza, were averaged over a 16-day window in May 2024.
+Table 1 summarizes the spatial responses distribution.
+The maximum inter-site distance among sites is approximately $42,909$ kilometers.
+Our central scientific objective is to jointly predict ndvi and rr from massive, globally distributed datasets of this scale using GeoAI.
+We assess whether multivariate statistical models can deliver accurate, scalable, and timely predictions across millions of spatial locations while accounting for dependence structures in the solar zenith angle.
+These capabilities are crucial for application in ecology, agriculture, and climate policy.
+5.2 Results using dbps
+To analyze our Vegetation Index (vi) data, we follow machine-generated exploratory data analysis (eda) described in Section LABEL:sec:eda.
 The predictors comprise an intercept and the solar zenith angle for that location ($p=2$).
-The latter depicts the strength of solar irradiation.
-All the data were collected over a 16-day window in May 2024 and then averaged.
-[…]
+Based on Section LABEL:sec:eda, we set $\alpha\in\{0.825,0.909\}$, and $\phi\in\{0.049,0.067\}$ respectively.
+We specify $\{\gamma,\Sigma\}$ in (2.5) using $m_{0}=0_{p\times q}$, $M_{0}=10\mathbb{I}_{p}$, $\Psi_{0}=\mathbb{I}_{q}$, $\nu_{0}=3$, and finally we opt for an exponential spatial covariance function.
+Following insights from Section LABEL:sec:sim5_sens, we fix the subset size at $n_{k}\in\{250,500\}$, leading to a number of subsets $K\in\{4,000\;,\;2,000\}$ respectively.
+We use a random scheme to form the partition of the dataset and present results when $K=2,000$ and in Table 2 for $K=4,000$.
 The multivariate model developed in Section 2.3 with ndvi and rr comprising the $q=2$ columns of $Y$.
-Figures 7 and 8 illustrate maps corresponding to ndvi and rr, respectively, using $K=2,000$ subsets.
+Figures 6 and 7 illustrate maps corresponding to ndvi and rr, respectively, using $K=2,000$ subsets.
 The top left panel in each figure presents the spatially interpolated map of the training data for the respective responses.
@@
 Results correspond to $K=2,000$.
-Figure 8: Left to right: Maps for training data (top left), test data (top right) and predicted surface (bottom right) for rr.
+Figure 7: Left to right: Maps for training data (top left), test data (top right) and predicted surface (bottom right) for $r$.
 Empirical coverage for held-out values of outcomes (bottom left).
 Results correspond to $K=2,000$.
-Figures 7 and 8 show results from double bps.
+Figures 6 and 7 show results from double bps.
 Maps of interpolated posterior means and predicted responses are displayed in the bottom right of the respective figures.
-The bps interpolated maps are almost indistinguishable from the true surface (top right).
-This suggests that the automated bps effectively, perhaps even strikingly, recovers the spatial patterns in spite of the modeling simplifications over more elaborate statistical models (Banerjee,, 2020; Zhang et al.,, 2021).
-Moreover, the double bps behaves well in terms of predictive performance and uncertainty quantification for both values of $K$ considered.
+The dbps interpolated maps are almost indistinguishable from the true surface (top right).
+This suggests that the automated dbps effectively, perhaps even strikingly, recovers the spatial patterns in spite of the modeling simplifications over more elaborate statistical models (Banerjee, , 2020; Zhang et al., , 2021).
+The predictive $95\%$ credible intervals when $K=2,000$ for both responses are shown in the bottom left of Figures 6 and 7, respectively.
+Table 2 compares model parameters posterior estimates for dbps with $K\in\{4,000\;,\;2,000\}$, and the Bayesian conjugate linear model, which does not take into account spatial variability.
+The notably higher magnitudes of the intercepts in the non-spatial linear model are unsurprising as the spatial random effects absorb much of the variation in the intercepts.
+The solar zenith angle is positively associated with rr and negatively associated with ndvi in the non-spatial model.
+This, too, is expected since higher levels of solar irradiation are associated with higher red reflectance and with more arid regions with less vegetation.
+However, the spatial models reveal that the slope for ndvi is significantly positive with solar irradiation after the spatial effects have absorbed previously unaccounted latent or lurking factors that might have contributed to the negative slopes in the non-spatial model.
+Parameter Conjugate Linear model dbps (K=4, [^0] K=4{,}000) dbps (K=2, [^0] K=2{,}000)
+$\beta_{0,\textsc{ndvi}}$ 34.495 (34.392, 34.601) 1.364 (-1.039, 3.907) 1.767 (-0.422, 3.917)
+$\beta_{1,\textsc{ndvi}}$ -2.708 (-2.719, -2.697) 0.744 (0.478, 0.998) 0.707 (0.485, 0.939)
+$\beta_{0,\textsc{rr}}$ -16.924 (-17.014, -16.837) -0.418 (-2.470, 1.710) -0.562 (-2.187, 0.962)
+$\beta_{1,\textsc{rr}}$ 2.664 (2.655, 2.674) 0.945 (0.726, 1.158) 0.959 (0.795, 1.139)
+$\Sigma_{\textsc{ndvi}}$ 0.221 (0.220, 0.221) 0.182 (0.130, 0.257) 0.119 (0.099, 0.146)
+$\Sigma_{\textsc{ndvi},\textsc{rr}}$ -0.167 (-0.168, -0.166) -0.114 (-0.173, -0.079) -0.076 (-0.094, -0.061)
+$\Sigma_{\textsc{rr}}$ 0.155 (0.154, 0.155) 0.115 (0.082, 0.170) 0.077 (0.062, 0.092)
+Table 2: Vegetation Index data analysis parameter estimates for candidate models. Parameter posterior summary 50 (2.5, 97.5) percentiles.
+Model Time (min) RMSPE $\varrho_{\textsc{ndvi},\textsc{rr}}$
+dbps (K=2, [^0] K=2{,}000) 81 [0.206, 0.159, 0.183] -0.844 (-0.880, -0.795)
+dbps (K=4, [^0] K=4{,}000) 21 [0.241, 0.197, 0.219] -0.816 (-0.874, -0.706)
+Conjugate Linear Model 12 [0.474, 0.404, 0.439] -0.921 (-0.928, -0.915)
+gbm 2 [0.419, 0.352, 0.385] -0.993 (–,–)
+drf 10 [0.420, 0.353, 0.386] -0.991 (–,–)
+dnn 26 [0.422, 0.354, 0.388] -0.987 (–,–)
+automl 44 [0.419, 0.352, 0.385] -0.993 (–,–)
+Table 3: Vegetation Index data analysis computing time in minutes, rmspe, and empirical correlation ($\varrho$) for candidate models.
+Root mean square prediction error(s) presentation [ndvi, red reflectance, average].
+In addition to double bps (dbps) and Bayesian multivariate linear regression, we expanded the analysis by including four competitive algorithms: distributed random forest (drf), gradient boosting (gbm), deep neural network (dnn), and a fully automatic machine learning algorithm (automl).
+As detailed in Section LABEL:sec:compdetails, we implement and perform the analysis using the h2o R package (Fryda et al., , 2024).
+As no method for multivariate models is available within the h2o framework, we fit these algorithms separately for ndvi, and rr, considering the same explanatory variables.
+Then, we compute the empirical correlation among the predictions.
+In Table 3, we offer a comparison of computational costs, predictive performances, and empirical correlation.
+The double bps dominates both running time and predictive performances.
 It achieves an rmspe of $(0.206,0.159)$ when $K=2,000$, and $(0.241,0.197)$ with $K=4,000$ subsets for ndvi and rr respectively.
 This is significantly lower (by almost 100%) than the standard Bayesian conjugate multivariate linear model, which produces rmspe of $(0.474,0.439)$ for the two vegetation indexes, respectively.
-The predictive $95\%$ credible intervals when $K=2,000$ for both responses are shown in the bottom left of Figures 7 and 8, respectively.
-[…]
-Conjugate Linear model bps ($K=4,000$) bps ($K=2,000$)
-$\beta_{0,\textsc{ndvi}}$ 34.495 ( 34.392, 34.601) 1.364 (-1.039, 3.907) 1.767 (-0.422, 3.917)
-$\beta_{1,\textsc{ndvi}}$2.708 ( -2.719, -2.697) 0.744 ( 0.478, 0.998) 0.707 ( 0.485, 0.939)
-$\beta_{0,\textsc{rr}}$16.924 (-17.014, -16.837) -0.418 (-2.470, 1.710) -0.562 (-2.187, 0.962)
-$\beta_{1,\textsc{rr}}$ 2.664 ( 2.655, 2.674) 0.945 ( 0.726, 1.158) 0.959 ( 0.795, 1.139)
-$\Sigma_{\textsc{ndvi}}$ 0.221 ( 0.220, 0.221) 0.182 ( 0.130, 0.257) 0.119 ( 0.099, 0.146)
-$\Sigma_{\textsc{ndvi},\textsc{rr}}$0.167 ( -0.168, -0.166) -0.114 (-0.173, -0.079) -0.076 (-0.094, -0.061)
-$\Sigma_{\textsc{rr}}$ 0.155 ( 0.154, 0.155) 0.115 ( 0.082, 0.170) 0.077 ( 0.062, 0.092)
-$\varrho_{\textsc{ndvi},\textsc{rr}}$0.921 ( -0.928, -0.915) -0.816 (-0.874, -0.706) -0.844 (-0.880, -0.795)
-rmspe [0.474, 0.404, 0.439] [0.241, 0.197, 0.219] [0.206, 0.159, 0.183]
-Time 12 21 81
-Table 3: Vegetation Index data analysis parameter estimates, rmspe, and computing time in minutes for candidate models.
-Parameter posterior summary 50 (2.5, 97.5) percentiles.
-Root mean square prediction error(s) presentation [ndvi, red reflectance, average].
-Recalling that all the analyses are produced on a personal laptop (with just 5 physical cores) with minimal human intervention, the total run time of only 21 minutes with $K=4,000$ and 81 minutes with $K=2,000$ are impressive confirm the quadratic dependence of the partition size discussed in Section LABEL:sec:data1.
+Actually, the double bps outperformed all the other competitive algorithms in terms of rmspe.
+This is attained at a fraction of the computational time for other methods, let alone double bps models multivariate distributions jointly, and offer full explainability.
+Estimates of $\varrho_{\textsc{ndvi}}$ and $\varrho_{\textsc{rr}}$ reveal a well-documented negative association between the two responses.
+Indeed, the spatial patterns in these indices are almost the reverse of each other as revealed in Figures 6 and 7.
+The conjugate Bayesian linear model estimates higher negative correlation, while dbps tends to underestimate it, considerably so as $n_{k}$ decreases.
+We attempted comparisons with the lmc and bart models, but omitted due they exceeded memory.
+Recalling that all the analyses are produced on a personal laptop (with just 5 physical cores) with minimal human intervention, the total run time of only 21 minutes with $K=4,000$ and 81 minutes with $K=2,000$ for double bps are impressive and confirm the quadratic dependence of the partition size discussed in Section LABEL:sec:sim5_sens.
 Moreover, the strong dependence on the number of $J$ competitive models assessed is worth noting.
-Here $J=4$ while in Section LABEL:sec:data1 $J=3$.
-This suggests a marginal computational burden for each competitive model.
+Here $J=4$, this suggests a marginal computational burden for each competitive model.
 6 Discussion
-This manuscript devises a statistical modeling component for an artificially intelligent geospatial system (GeoAI) tailored to analyze massive datasets with minimal human intervention.
-The contribution harnesses analytically accessible multivariate statistical distributions in conjunction with Bayesian predictive stacking to deliver rapid inference on multivariate processes and circumventing iterative algorithms that require, often extensive, human tuning or control.
-Our proposed GeoSpatial AI system is capable of handling truly massive amounts of spatial data that preclude accessibility even on high performance CPU architectures and we develop Bayesian transfer learning using double Bayesian predictive stacking.
-A few additional remarks are warranted.
+This manuscript devises a statistical modeling component for an artificially intelligent geospatial system (GeoAI) tailored to analyze massive data with minimal human intervention.
+The contribution harnesses analytically accessible multivariate statistical distributions in conjunction with Bayesian predictive stacking to deliver rapid inference by circumventing iterative algorithms that require, often extensive, human tuning.
+Our proposed GeoSpatial AI system relies on Bayesian transfer learning using double Bayesian predictive stacking to process massive amounts of streaming spatial data on high-performance CPU architectures.
+Some additional remarks are warranted.
 The development here has been elucidated with a hierarchical matrix-variate spatial process framework.
-While modeling simplifications have been introduced to minimize human intervention, we wish to emphasize that the bps seamlessly applies to more versatile but analytically intractable models.
-First, for spatially misaligned multivariate data where not all the variables may have been observed over the same set of locations, we encounter missing entries in $Y$ but are still able to deliver analytically tractable closed-form distribution theory (Zhang and Banerjee,, 2022).
-Second, should users wish to explore more complex models with unknown $\{\alpha,\phi\}$ or work with alternate multivariate models, then we can implement full Bayesian inference (now using MCMC, Variational Bayes or Laplace approximations) on each subset and invoke bps to pool across the subsets of the data.
-Time for analyzing each subset increase due to model complexity, but scaling inference to massive datasets is still enabled by bps—and that too with considerable efficiency over alternate divide-and-conquer approaches such as the geometric median, or gm, of posteriors (used in meta-kriging developed by Guhaniyogi and Banerjee,, 2018).
-A preliminary examination of Section 2.3 reveals that posterior inference is conceptualized as a “mixture of mixtures”.
+Although modeling simplifications have been introduced to minimize human intervention, we emphasize that dbps seamlessly applies to more versatile but analytically intractable models.
+For spatially misaligned multivariate data, where not all variables have been observed in the same set of locations, we encounter missing entries in $Y$ using analytically tractable closed-form distribution theory (Zhang and Banerjee, , 2022).
+Should users wish to explore more complex models with unknown $\{\alpha,\phi\}$ or work with alternate multivariate models, we can implement Bayesian inference (using MCMC, Variational Bayes or Laplace approximations) on each subset and invoke bps to pool across the subsets of the data.
+The time for analyzing each subset increases due to model complexity, but scaling inference to massive datasets is still achieved by bps—and this too with considerable efficiency over divide-and-conquer approaches such as the geometric median, or gm, of posteriors (see meta-kriging in Guhaniyogi and Banerjee, , 2018).
+Future research can build upon recent work by Cabel et al., (2025) to enrich spatial-temporal dependence structures within dbps and to further accelerate amortized inference.
+Section 2.3 devised posterior inference as a “mixture of mixtures”.
 Following Yao et al., (2018), we prefer stacking to fitting a mixture model because the former is numerically robust and requires almost no human tuning.
@@
 Disseminating our proposed product with our accompanying software, which is currently being migrated to R, is expected to significantly boost Geospatial AI systems.
-Future directions will also explore perceived potential of bps as a feeder for emerging amortized inference methods (Ganguly et al.,, 2023; Zammit-Mangion et al.,, 2024; Sainsbury-Dale et al.,, 2024) to achieve Bayesian inference.
-Rapid delivery of posterior estimates of the entire spatial process from bps will amount to more training data for amortized neural learners that can result in accelerated tuning for subsequent Bayesian inference.
+Future directions will also explore perceived potential of dbps as a feeder for emerging amortized inference methods (Ganguly et al., , 2023; Zammit-Mangion et al., , 2024; Sainsbury-Dale et al., , 2024) to achieve Bayesian inference.
+Rapid delivery of posterior estimates of the entire spatial process from dbps will amount to more training data for amortized neural learners that can result in accelerated tuning for subsequent Bayesian inference.
 We do not see our proposed approach as a competitor to, but rather as supplementary to amortized neural inference.
+Refer to Section 4.2 for an effective representation.
 Such developments will be pursued as future research.
@@
 Supplementary Materials
-The online supplement includes additional derivations and details on matrix-variate distributions (Section LABEL:sec:distributiontheory), computational details and algorithms (Section LABEL:sec:compdetails), additional simulation experiment results (Section LABEL:sec:sim_suppl), details on accelerated bps for univariate spatial data (Section LABEL:sec:AccelerUnivarMod) and analysis of an additional dataset on sea surface temperature (Section LABEL:sec:data1).
-Section LABEL:sec:eda outlines automated computation of model hyperparameters without human intervention.
-Computer programs to reproduce all experiments and data analysis are publicly accessible from the GitHub repository lucapresicce/Bayesian-Transfer-Learning-for-GeoAI
+The online supplement includes exploratory data analyses, and automated settings for model hyperparameters without human intervention (Section LABEL:sec:eda), additional derivations and details on matrix-variate distributions (Section LABEL:sec:distributiontheory), some theoretical insights (Section LABEL:sec:upperbound), computational details and algorithms (Section LABEL:sec:compdetails), and additional simulation experiment results (Section LABEL:sec:sim_suppl_M) Computer programs to reproduce all our analysis are publicly accessible from the GitHub repository lucapresicce/Bayesian-Transfer-Learning-for-GeoAI